### v2

# Cancellation in sums over special sequences on $\mathbf{{\rm{GL}}_{m}}$ and their applications

Abstract.
Let $a(n)$ be the $n$-th Dirichlet coefficient of the automorphic $L$-function or the Rankin–Selberg $L$-function. We investigate the cancellation of $a(n)$ over specific sequences linked to the Waring–Goldbach problem, by establishing a nontrivial bound for the additive twisted sums over primes on ${\mathrm{GL}}_{m}$ and ${\mathrm{GL}}_{m}\times{\mathrm{GL}}_{m}.$ The bound does not depend on the generalized Ramanujan conjecture or the nonexistence of Landau–Siegel zeros. Furthermore, we present an application associated with the Sato–Tate conjecture and propose a conjecture about Goldbach’s conjecture on average bound.
Key words and phrases:
Exponential estimates, Automorphic representations, Waring–Goldbach problem
2020 Mathematics Subject Classification:
11L07, 11F66, 11P55
1. introduction
In the field of number theory, the additive and multiplicative structures of integers are two major research subjects. Multiplicative number theory investigates the distribution of prime numbers through the analysis of generating functions, with the Prime Number Theorem and Dirichlet’s theorem being the most famous classical achievements in this field. On the other hand, additive number theory focuses on exploring the representation of additive decomposition of integers, with the Waring–Goldbach problem being a well-known subject in this field. It has been generally assumed that these two structures are independent; however, so far there has been no proof to confirm this assumption. Problems involving both structures tend to be more challenging to solve, such as the twin prime conjecture, which asserts that there should be infinitely many pairs of primes which differ by 22 22. Unfortunately this conjecture remains open, but recently Zhang [42], Maynard [27], Polymath [30] and many other scholars have made significant progress on it. Merging both additive and multiplicative aspects within number theory, this paper specifically explores coefficients of the standard $L$-functions over sequences related to the Waring–Goldbach problem, where the Dirichlet coefficients of $L$-functions are the generating functions.
Let $\mathcal{P}$ be the set of all primes, $N$ be a sufficiently large positive integer. Define the sets
$\mathcal{P}_{k}(N):=\{p\in\mathcal{P},p^{k}\leq N\},\quad$
and
$\displaystyle\mathcal{J}_{k,u}(N):=$$\displaystyle\{(p_{1},p_{2},\dots,p_{u})|\ p_{i}\in\mathcal{P}_{k}(N)~{}(1\leq i\leq u),N=p_{1}^{k}+p_{2}^{k}+\cdots+p_{u}^{k}\},$
where $k$ and $u$ are positive integers. For a fixed $k$, finding the smallest $u$ such that the set $\mathcal{J}_{k,u}(N)$ is non-empty and providing the asymptotic formula for the number of elements in $\mathcal{J}_{k,u}(N)$, are the objectives of the Waring–Goldbach problem. We introduce two special cases associated with $\mathcal{J}_{k,u}(N)$: the linear case $k=1$ and the quadratic case $k=2$. In both cases, $N$ is an odd integer.
For $k=1$, Vinogradov [39] proved
#$\#\mathcal{J}_{1,3}(N)=\frac{1}{2}(1+o(1))\mathfrak{G}_{1,3}(N)\frac{N^{2}}{(\log N)^{3}},$
where $\mathfrak{G}_{1,3}(N)$ is the singular series
$\mathfrak{G}_{1,3}(N)=\prod_{p\mid N}\left(1-\frac{1}{(p-1)^{2}}\right)\prod_{p\nmid N}\left(1+\frac{1}{(p-1)^{3}}\right).$ (1.1)
The equation implies that every sufficiently large odd integer $N$ can be expressed as the sum of three primes, representing Vinogradov’s three primes theorem. Furthermore, $u=3$ is the minimum number of variables required for $\mathcal{J}_{1,u}(N)$ to be non-empty. For the quadratic case $k=2$, Hua [16] gave the best result and provided
#$\#\mathcal{J}_{2,5}(N)>0\quad\text{when}\ N\equiv 5\ ({\rm mod}\ 24).$
The number of variables required is still the best ( more details see Lemma 7.1).
Inspired by the above result, many scholars consider the analogue of Vinogradov’s three primes theorem for the coefficients of the standard $L$-functions. Let $c$ be a positive number, but its value may differ between instances throughout the paper. Fouvry and Ganguly [8] proved that there exists an effective constant $c>0$ such that
$\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\nu_{f}(p_{1})\ll_{f}N^{2}\exp\left(-c\sqrt{\log N}\right),$ (1.2)
where $\nu_{f}(n)$ is the normalized Fourier coefficient of a primitive holomorphic or Maass cusp form $f$. Let $F$ be a Hecke–Maass form for $\mathrm{SL}(3,\mathbb{Z})$, with $A_{F}(n,1)$ denoting the $n$th coefficient of the Dirichlet series for it. Hou, Jiang and Lü [15] proved that there exists an effective constant $c>0$ such that
$\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}A_{F}(p_{1},1)\ll_{F}N^{2}\exp(-c\sqrt{\log N}).$ (1.3)
For a primitive Dirichlet character $\chi$, Hoffstein and Ramakrishnan [14, Theorem A] and Banks [1] have proved that the twisted $L$-functions $L(s,f\otimes\chi)$ and $L(s,F\otimes\chi)$ don’t exist Landau–Siegel zeros. The above two estimates depend on the nonexistence of Landau–Siegel zeros. Since the existence of Landau–Siegel zeros is still open for the $L$-functions with higher ranks, the method of the above two estimates can not be used to settle the similar problems for higher ranks. In this paper, we apply a new approach not to only obtain the relating estimates for $L$-functions and Rankin–Selberg $L$-functions with higher ranks but also obtain the same estimates (1.2) and (1.3). Moreover, our method can be applied to consider the cancellation of the coefficients associated with the quadratic case. The method in quadratic case is more ingenious. To obtain the optimal upper bound, we divide a subinterval of $[0,1)$ twice to select the appropriate exponential estimates (more details see Section 5.3). It is a common approach to handle the Waring–Goldbach problem.
Let $\mathbb{A}$ be the ring of adeles over $\mathbb{Q}$. Define a unitary Hecke character $\omega:\mathbb{Q}^{\times}\backslash\mathbb{A}^{\times}\rightarrow\mathbb{C}^{\times}$. Consider the set
$\mathfrak{E}_{m}:=\{\pi\subset L^{2}(\mathrm{GL}_{m}(\mathbb{Q})\backslash\mathrm{GL}_{m}(\mathbb{A});\omega)\mid\pi\text{ is irreducible and cuspidal}\}.$
Given $\pi\in\mathfrak{E}_{m}$, let $\mathfrak{q}_{\pi}$ be the conductor of $\pi$, $L(s,\pi)$ be the associated standard $L$-function, and $\widetilde{\pi}\in\mathfrak{E}_{m}$ be the contragredient representation. We write $\lambda_{\pi}(n)$ to be the $n$-th Dirichlet coefficient of $L(s,\pi)$. As is well known, $\chi\in\mathfrak{E}_{1}$ corresponds to a Hecke character of the idele class group ×\$\mathbb{Q}^{\times}\backslash\mathbb{A}^{\times}$. We define the set
$\mathfrak{E}_{m}^{b}:=\{\pi\in\mathfrak{E}_{m}$ $|\pi=\widetilde{\pi}\text{ and }\pi\neq\pi\otimes\chi$
[…]
We will examine the cancellation of the multiplicative Dirichlet coefficients $\lambda_{\pi}(n)$ over specific sequences $\mathcal{J}_{1,3}(N)$ and $\mathcal{J}_{2,5}(N)$, and give some applications.
Theorem 1.1.
Let $\pi\in\mathfrak{E}_{m}^{b}$ and $B>0$ be a sufficiently large constant. Then
$r_{\pi}(N):=\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\lambda_{\pi}\left(p_{1}\right)\ll_{\pi,B}N^{2}(\log N)^{-B},$
and
$w_{\pi}(N):=\sum_{(p_{1},p_{2},p_{3},p_{4},p_{5})\in\mathcal{J}_{2,5}(N)}\lambda_{\pi}\left(p_{1}\right)\ll_{\pi,B}N^{3/2}(\log N)^{-B}.$
The bounds in the above theorem are weaker than the results (1.2) and (1.3). This is because we consider the above result under the existence of Landau–Siegel zeros for the $L$-functions with higher ranks. The zero-free region used in the above theorem restricts to the condition $\pi\in\mathfrak{E}_{m}^{b}$. Under a hypothesis concerning an enhanced zero-free region for $\pi\in\mathfrak{E}_{m}$, we can achieve better upper bounds for more general $L$-functions.
Hypothesis S.
Let $1\displaystyle 1/2\leq\delta<$ 1. For any $\pi\in\mathfrak{E}_{m}$ with $m\geq 1$, the $L$-functions $L(s,\pi)$ have no zeros for $\mathrm{Re}(s)>\delta$.
Theorem 1.2.
Let $\pi\in\mathfrak{E}_{m}$ and $\varepsilon$ be a sufficiently small positive number. Under Hypothesis S,
$r_{\pi}(N)\ll_{\varepsilon}m^{2}N^{\frac{5}{3}+\frac{\delta}{3}+\varepsilon}(\log(N\mathfrak{q}_{\pi}))^{3},$
and
$w_{\pi}(N)\ll_{\varepsilon}m^{2}(N^{\frac{23}{16}+\varepsilon}+N^{\frac{4}{3}+\frac{\delta}{6}+\varepsilon})(\log(N\mathfrak{q}_{\pi}))^{3}.$
Remark.
  1. (1)
    Regarding the upper bounds in (1.2) and (1.3), we can use the same method in Theorem 1.2 to obtain them.
  2. (2)
    Note that the upper bound of $w_{\pi}(N)$ consists of two terms. This is because we use the Dirichlet approximation theorem to divide the minor arcs twice. Such an approach is effective in handling minor arcs within the circle method while enlarging major arcs.
In our above results, we have no asymptotic formulas. But the upper bounds we obtain are better than the error terms in the classical cases in Lemma 7.1, which is a reflection of the oscillation property of $\lambda_{\pi}(n)$ in the relating additive structure. Furthermore, we want to obtain asymptotic formulas concerning the Rankin–Selberg coefficients over $\mathcal{J}_{k,u}(N)$. Let us first fix some notation. For $\mathrm{Re}(s)>1$,
$-\frac{L^{\prime}}{L}(s,\pi\times\tilde{\pi})=\sum_{n=1}^{\infty}\frac{\Lambda(n)a_{\pi\times\tilde{\pi}}(n)}{n^{s}}$
where $\Lambda(n)$ is the von Mangoldt function defined by
$\Lambda(n):=\begin{cases}\log p,&\text{ if }n=p^{k},\\ 0,&\text{ otherwise},\end{cases}$
and
$a_{\pi\times\tilde{\pi}}(n)=\begin{cases}\sum_{j=1}^{m^{2}}\alpha_{\pi\times\tilde{\pi},j}(p)^{k},&\text{if }n=p^{k},\\[6.0pt] 0,&\text{otherwise},\end{cases}$
for all primes $p$ and integers $k\geq 1$. For $j=1,2,\ldots,m^{2}$, $\alpha_{\pi\times\tilde{\pi},j}(p)$ are given below after (3.9), which are local parameters associated with $\pi\times\tilde{\pi}$. We give a milder hypothesis than generalized Ramanujan conjecture (see Section 33 33) associated to Rankin–Selberg $L$-functions. In a number of special situations, the following hypothesis may be dropped, and we provide a few examples after Theorem 1.3. Let $x$ be a sufficiently positive, large number.
Hypothesis A.
Let $\pi\in\mathfrak{E}_{m}$, then
$\sum_{n\leq x}\Lambda(n)a_{\pi\times\tilde{\pi}}(n)^{2}\ll_{\pi}x.$
Under the above hypothesis and the existence of Landau–Siegel zeros for Rankin–Selberg $L$-functions, we obtain the following result with a new zero-free region for $L$-functions in [11].
Theorem 1.3.
Let $\pi\in\mathfrak{E}_{m}$, $\varepsilon$ be a sufficiently small positive number and $B>0$ be a sufficiently large constant. Under Hypothesis A, we get
$r_{\pi\times\tilde{\pi}}(N):=\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\lambda_{\pi\times\tilde{\pi}}\left(p_{1}\right)=\#\mathcal{J}_{1,3}(N)+O_{\pi,B}(N^{2}(\log N)^{-B}),$
and
$w_{\pi\times\tilde{\pi}}(N):=\sum_{(p_{1},p_{2},p_{3},p_{4},p_{5})\in\mathcal{J}_{2,5}(N)}\lambda_{\pi\times\tilde{\pi}}\left(p_{1}\right)=\#\mathcal{J}_{2,5}(N)+O_{\pi,B}(N^{2}(\log N)^{-B}).$
Remark.
Following from Lemma 7.1 in the Appendix, we note that if $N$ satisfies some necessary congruence conditions, then #$\#\mathcal{J}_{1,3}(N)\asymp N^{2}/(\log N)^{3}$ and #$\#\mathcal{J}_{2,5}(N)\asymp N^{3/2}/(\log N)^{5}$ which are the main terms in the above asymptotic formulas.
Example 1.
If $\pi\in\mathfrak{E}_{m}$ satisfies the generalized Ramanujan conjecture, Hypothesis A in Theorem 1.3 can be easily dropped. We will give an explicit example. For a positive integer $N_{1}$ and an even positive integer $k_{1}$, we denote by $H_{k_{1}}^{*}(N_{1})$ the finite set of all primitive cusp forms of weight $k_{1}$ with trivial nebentypus for the Hecke congruence group $\Gamma_{0}(N_{1})$. Each $f\in H_{k_{1}}^{*}(N_{1})$ has a Fourier expansion
$f(z)=\sum_{n=1}^{\infty}\lambda_{f}(n)n^{\frac{k_{1}-1}{2}}e(nz)$
in the upper half-plane. It is known that Newton and Thorne [28, 29] have proven that $\mathrm{sym}^{j}f$ corresponds to $\mathrm{sym}^{j}\pi_{f}$, which is a cuspidal automorphic representation of $\mathrm{GL}_{j+1}$ over $\mathbb{Q}$ for all $j\geq 1$. In this scenario, the representations $\mathrm{sym}^{j}\pi_{f}~{}(j\geq 1)$ satisfy the generalized Ramanujan conjecture (see Section 6). Thus, we obtain
$\sum_{n\leq x}\Lambda(n)a_{\mathrm{sym}^{j}\pi_{f}\times\widetilde{\mathrm{sym}^{j}\pi_{f}}}(n)^{2}\ll_{\pi_{f},j}x,\quad j\geq 1.$
Example 2.
If $\pi\in\mathfrak{E}_{2}$, we use the relationship between Rankin–Selberg $L$-functions and adjoint $L$-functions to drop Hypothesis A. The construction of an adjoint $L$-function is described by employing a Rankin–Selberg $L$-function and zeta function:
$L(s,\operatorname{Ad}\pi):=\frac{L(s,\pi\times\tilde{\pi})}{\zeta(s)},$
where $\operatorname{Ad}\pi$ is the adjoint lift of $\pi$. The adjoint $L$-function is meromorphic since $L(s,\pi\times\widetilde{\pi})$ and $\zeta(s)$ are meromorphic. Details regarding the Rankin–Selberg $L$-functions can be found in Section 3. For $\mathrm{Re}(s)>0$, we express
$L(s,\operatorname{Ad}\pi)=\prod_{p}\prod_{i=1}^{m}\prod_{\begin{subarray}{c}j=1\\ i\neq j\end{subarray}}^{m}\left(1-\frac{\alpha_{i,\pi}(p)\overline{\alpha_{j,\pi}(p)}}{p^{s}}\right)^{-1}=\sum_{n=1}^{\infty}\frac{\lambda_{\operatorname{Ad}\pi}(n)}{n^{s}}.$
Taking the logarithmic derivative for $L(s,\operatorname{Ad}\pi)$, we obtain that, for $\mathrm{Re}(s)>1$,
$\Lambda(n)a_{\pi\times\tilde{\pi}}(n)=\Lambda(n)(1+a_{\operatorname{Ad}\pi}(n)).$ (1.4)
If $m=2$, the adjoint lift of $\pi$ is known to be automorphic. Furthermore, $\operatorname{Ad}\pi$ is cuspidal if and only if $\pi$ is not an automorphic induction of a Hecke character according to [10]. Thus, by (3.15) and (1.4), we obtain that
$\sum_{n\leq x}\Lambda(n)a_{\pi\times\tilde{\pi}}(n)^{2}\ll_{\pi}x.$
Example 3.
If $\pi\in\mathfrak{E}_{2}$ and $\mathrm{sym}^{2}\pi$ represents the second symmetric power of the standard representation of $\mathrm{GL}_{2}$, Hypothesis A for the special case $\mathrm{sym}^{2}\pi$ in Theorem 1.3 can be dropped. The automorphy of $\mathrm{sym}^{j}\pi$ was established for $j=2$ by Gelbart and Jacquet [10], and for $j=3,4$ by Kim and Shahidi [24, 25]. Following from well-known identities: for all $j\geq 1$, the equations
$\mathrm{sym}^{2}\left(\mathrm{sym}^{j}\pi\right)=\bigoplus_{0\leq i\leq j/2}\mathrm{sym}^{2j-4i}\pi$
(see [9, 159]) and
$\left(\mathrm{sym}^{j}\pi\right)^{\otimes 2}:=\mathrm{sym}^{j}\pi\otimes\mathrm{sym}^{j}\pi=\bigoplus_{0\leq r\leq j}\mathrm{sym}^{2r}\pi$ (1.5)
(see [9, 151]) where $\mathrm{sym}^{0}\pi_{f}$ is the 1 -dimensional trivial representation, we have that
$\operatorname{Ad}\mathrm{sym}^{j}\pi=\bigoplus_{1\leq r\leq j}\mathrm{sym}^{2r}\pi.$
Thus, $\operatorname{Ad}\mathrm{sym}^{2}\pi$ is cuspidal and automorphic. If $\pi\in\mathfrak{E}_{2}$, by (3.15) and (1.4), we obtain that
$\sum_{n\leq x}\Lambda(n)a_{\mathrm{sym}^{2}\pi\times\widetilde{\mathrm{sym}^{2}\pi}}(n)^{2}\ll_{\pi}x.$
As an application, we present an analogue of the Sato–Tate conjecture. We restrict our analysis to forms with trivial nebentypus, ensuring that all Fourier coefficients are real. Additionally, we assume that $f\in H_{k_{1}}^{*}(N_{1})$ does not possess complex multiplication (CM), meaning there is no imaginary quadratic field $K$ such that for $p\nmid N_{1}$, $p$ is inert in $K$ if and only if $\lambda_{f}(p)=0$. Deligne’s proof [7] of the Weil conjectures implies that for each prime $p$, there exists an angle $\theta_{p}\in[0,\pi]$ such that $\lambda_{f}(p)=2\cos\theta_{p}$. The Sato–Tate conjecture asserts that if $f\in H_{k_{1}}^{*}(N_{1})$ is non-CM, then the sequence $\left\{\theta_{p}\right\}$ is equidistributed in the interval $[0,\pi]$ with respect to the measure $\mathrm{d}\mu_{\mathrm{ST}}:=(2/\pi)\sin^{2}\theta\mathrm{d}\theta$. Equivalently, we obtain that
$\pi_{f,I}(x):=\#\left\{p\leq x:\theta_{p}\in I,p\nmid N_{1}\right\}\sim\mu_{\mathrm{ST}}(I)\pi(x)\quad\text{ as }x\rightarrow\infty,$ (1.6)
where $\pi(x)=\#\{p\leq x\}$ and $I=[\alpha,\beta]\subseteq[0,\pi]$. Barnet-Lamb, Geraghty, Harris, and Taylor [2] proved the Sato–Tate conjecture. For all primitive characters $\chi$, Thorner [37] has proved that the twisted $L$-functions $L(s,\mathrm{sym}^{j}f\otimes\chi)$ have no Landau–Siegle zeros. In this paper, we give a simplier proof in Lemma 6.1. Based on this property, we will show that the equidistribution of $\theta_{p_{1}}$ remains undisturbed over the set $\mathcal{J}_{1,3}(N)$. Moreover, as an application of Theorem 1.2, we will also investigate the distribution of $\lambda_{f}(p_{1})$ associated with $(p_{1},p_{2},p_{3})$ in the set $\mathcal{J}_{1,3}(N)$ under Hypothesis S. We define
$\mathcal{J}_{1,3,f,I}(N):=\{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N):\theta_{p_{1}}\in I,p_{1}\nmid N_{1}\}.$
Theorem 1.4.
Fix a non-CM form $f\in H_{k_{1}}^{*}(N_{1})$.
  1. (1)
    We have
    #$\#\mathcal{J}_{1,3,f,I}(N)=\mu_{\mathrm{ST}}(I)\left(\#\mathcal{J}_{1,3}(N)\right)+O\left(\frac{\log(k_{1}N_{1}\log N)}{(\log N)^{1/2}}\left(\#\mathcal{J}_{1,3}(N)\right)\right).$
  2. (2)
    Under Hypothesis S, we have
    #$\#\mathcal{J}_{1,3,f,I}(N)=\mu_{\mathrm{ST}}(I)\left(\#\mathcal{J}_{1,3}(N)\right)+O(N^{11/6+\delta/6}(\log N)^{3}).$
It follows from the above result and (1.6) that
#$\frac{\#\mathcal{J}_{1,3,f,I}(N)}{\#\mathcal{J}_{1,3}(N)}\sim\frac{\pi_{f,I}(N)}{\pi(N)}\quad\text{ as }N\rightarrow\infty.$ (1.7)
From a different perspective, based on (1.7) and Lemma 7.1, then
#$\frac{\#\mathcal{J}_{1,3,f,I}(N)}{\pi_{f,I}(N)}\sim\frac{\#\mathcal{J}_{1,3}(N)}{\pi(N)}\sim\frac{\mathfrak{G}_{1,3}(N)N}{2(\log N)^{2}}\quad\text{ as }N\rightarrow\infty.$ (1.8)
The Goldbach conjecture states that every even number $2N$ can be expressed as the sum of two odd primes. Hardy and Littlewood [12] conjectured an asymptotic formula for the number of such representations as follow:
#$\#\mathcal{J}_{1,2}(2N)\sim 2C_{2}\frac{2N}{(\log 2N)^{2}}\prod_{\mathfrak{p}}\left(\frac{\mathfrak{p}-1}{\mathfrak{p}-2}\right),$ (1.9)
where $\mathfrak{p}$ is an odd prime divisor of $2N$, and
$C_{2}=\prod_{\varpi=3}^{\infty}\left(1-\frac{1}{(\varpi-1)^{2}}\right).$
Comparing the result (1.8) with the conjecture (1.9), it seems that the additional selection condition of $p_{1}$ in $\mathcal{J}_{1,3}(N)$ does not affect the selection of $p_{2}$ and $p_{3}$ in $\mathcal{J}_{1,3}(N)$. Following from some results on the Waring–Goldbach problem, it is observed that any $p_{i}$ ($1\displaystyle 1\leq i\leq k-$ 1), existing in $\mathcal{J}_{k,u}(N)$, covers almost all primes. Therefore, we aim to consider Goldbach’s conjecture from a probabilistic perspective.
Define the set
[…]
$\displaystyle(2,2,N-4),...,(2,N-3,1),\dots,(N-3,1,2),(N-3,2,1),(N-2,1,1)\}.$
Let $\mathcal{F}$ be the set of all subsets of $\Omega$ and $\phi$ be an empty set. According to [41, Chapter 1], $(\Omega,\mathcal{F})$ forms a measurable space. Let $\mathbf{P}$ be a map: $\mathcal{F}\rightarrow[0,1]$, which is a probability measure on $(\Omega,\mathcal{F})$ satisfying
$\left\{\begin{array}[]{c}\mathbf{P}(\phi)=0,\quad\mathbf{P}(\Omega)=1,\\ A_{i}\in\mathcal{F}~{}(i=1,2,\dots,2^{\#\Omega}),\quad\mathbf{P}\left(A_{i}\right)=\frac{\#A_{i}}{\#\Omega}.\end{array}\right.$
The triple $(\Omega,\mathcal{F},\mathbf{P})$ forms a probability space. We denote two sets
$A:=\{(p_{1},n_{2},n_{3})|p_{1}\in\mathcal{P},n_{2},n_{3}\in\mathbb{N^{+}},N=p_{1}+n_{2}+n_{3}\}$
and
$A^{\prime}:=\{(n_{1},p_{2},p_{3})|n_{1}\in\mathbb{N^{+}},p_{2},p_{3}\in\mathcal{P},N=n_{1}+p_{2}+p_{3}\}.$
It is evident that $A\cap A^{\prime}$ is the set $\mathcal{J}_{1,3}(N)$. Inspired by (1.8), we propose an interesting conjecture related to Goldbach’s conjecture.
Conjeture 1.5.
As for $(\Omega,\mathcal{F},\mathbf{P})$, the events $A$ and $A^{\prime}$ satisfy
$\mathbf{P}(A\cap A^{\prime})\sim\mathbf{P}(A)\mathbf{P}(A^{\prime}).$
By the Prime Number Theorem and Lemma 7.1, when $N$ is a sufficiently large odd integer, we derive that
$\mathbf{P}(A)\sim\frac{1}{\log N},\quad\quad\mathbf{P}(A\cap A^{\prime})\sim\frac{\mathfrak{G}_{1,3}(N)}{(\log N)^{3}}.$
Under Conjecture 1.5,
$\mathbf{P}(A^{\prime})\sim\frac{\mathbf{P}(A\cap A^{\prime})}{\mathbf{P}(A)}\sim\frac{\mathfrak{G}_{1,3}(N)}{(\log N)^{2}}.$
And since the number of samples in $\Omega$ is $(N^{2}-3N+2)/2$, then
#$\#A^{\prime}\sim\frac{\mathfrak{G}_{1,3}(N)N^{2}}{2(\log N)^{2}}.$ (1.10)
Also, we calculate the number of the samples in $A^{\prime}$, that is,
#$\#\mathcal{J}_{1,2}(N-4)+\#\mathcal{J}_{1,2}(N-5)+\dots+\#\mathcal{J}_{1,2}(4).$ (1.11)
It follows from (1.10) and a rough estimate that
$\sum_{n=2}^{(N-5)/2}\#\mathcal{J}_{1,2}(2n)\sim\frac{\mathfrak{G}_{1,3}(N)N^{2}}{2(\log N)^{2}}.$ (1.12)
If Conjecture 1.5 holds, we would obtain the estimate (1.12), which shows that Goldbach’s conjecture is right on average.
Remark.
For (1.11), we can use (1.9) to obtain a precise estimate
$\sum_{n=2}^{(N-5)/2}2C_{2}\frac{2n}{(\log 2n)^{2}}\prod_{\mathfrak{p}}\left(\frac{\mathfrak{p}-1}{\mathfrak{p}-2}\right).$ (1.13)
Comparing (1.13) with (1.12), Conjecture 1.5 can also provide a basis for the correctness of the asymptotic formula (1.9).
Structure of this paper. The structure of this paper is outlined as follows. In Section 22 22, we give three important propositions, which are used to prove our main theorems. we give some fundamental definitions and properties of $L$-functions along with some useful lemmas in Section 3. Section 4 is dedicated to the proof of Propositions 2.12.3, leveraging the Rankin–Selberg results, the Mellin transform, and functional equations of $L$-functions. Proceeding to Section 5, we establish Theorems 1.11.3 by employing the Hardy–Littlewood circle method in conjunction with Propositions 2.12.3. In Section 6, we obtain that the twisted $L$-functions $L(s,\mathrm{sym}^{j}f\otimes\chi)$ have no Landau-Siegle zeros. And an analogue of the Sato–Tate conjecture is presented as an application of method in Theorem 1.2, denoted as Theorem 1.4. Finally, Section 7 provides several useful appendices.
In this paper, we investigate the apparent oscillation properties of $\lambda_{\pi}(n)$ in interactions with additive structure of integers. For example, in Theorem 1.4, we observe that the equidistribution of the sequence $\theta_{p}$ in the interval $[0,\pi]$ is not disturbed by the additive structure of intergers associated with the Waring–Goldbach problem. We hope that the ideas in this paper (particularly the combination of ideas from both additive problem and multiplicative problems) may inspire further applications and advancements in related research areas.
2. Exponential estimates
In relation to the Waring–Goldbach problem, we will utilize the Hardy–Littlewood circle method. The idea of this method involves transforming a discrete weighted sum into an integral over $[0,1)$ with respect to related exponential sums such as
$\sum_{n\leq x}\Lambda(n)a(n)e\left(n^{k}\alpha\right),$ (2.1)
where $a(n)$ is an arithmetic function and $\alpha\in[0,1)$. Estimating exponential sums is the key to solve the problem. For any $\alpha\in[0,1)$, Fouvry and Ganguly [8] obtained
$\sum_{n\leq x}\Lambda(n)\nu_{f}(n)e(n\alpha)\ll_{f}x\exp(-c\sqrt{\log x}).$ (2.2)
Hou, Jiang and Lü [15] proved
$\sum_{n\leq x}\Lambda(n)A_{F}(n,1)e(n\alpha)\ll_{F}x\exp(-c\sqrt{\log x}).$ (2.3)
In the two equations above, $c$ is an effective constant. The two upper bounds above were used to get (1.2) and (1.3), respectively. The nonexistence of Landau–Siegel zeros plays an important role in estimating the sums (2.2) and (2.3). Under the assumption of the nonexistence of Landau–Siegel zeros and some other assumptions, Jiang and Lü [21] considered (2.1), with $k\geq 1$ and $a(n)$ being the Fourier coefficient of a Hecke–Maass form for $\mathrm{SL}(m,\mathbb{Z})$, $m\geq 4$.
In this paper, depending not on the nonexistence of Landau–Siegel zeros in the twisted $L$-functions, we consider the sum (2.1) for high ranks. Let $\mathbb{T}$ be the torus $\\\displaystyle\mathbb{Z}\backslash\mathbb{R}$, and identify $\mathbb{T}$ with the unit interval $[0,1)$ throughout this paper. For any $\alpha\in\mathbb{T}$, we will provide some estimates of (2.1) with $a(n)$ being $\lambda_{\pi}(n)$ or $\lambda_{\pi\times\tilde{\pi}}(n),$ which do not depend on the generalized Ramanujan conjecture (GRC in brief, described specifically in Section 3) or the nonexistence of Landau–Siegel zeros in the twisted $L$-functions for the first time.
Proposition 2.1.
Let $\pi\in\mathfrak{E}_{m}^{b}$. For any real number $\alpha\in\mathbb{T}$, suppose that there exist two integers $a$ and $q$, such that
[…]
where $Q$ is a large positive integer. For all $A>0$, if $q\leq(\log x)^{A}$, then
$\sum_{p\leq x}\lambda_{\pi}(p)e(p^{k}\alpha)\log p\ll_{\pi,A}\bigg{(}1+\frac{x^{k}}{qQ}\bigg{)}q^{\frac{m+1}{2}}x\exp\left(-c\sqrt{\log x}\right),$
where $c$ is an ineffective constant depending on $A$ and $\pi$.
Proposition 2.2.
Let $\pi\in\mathfrak{E}_{m}$, $\varepsilon$ be a sufficiently small positive number. For any real number $\alpha\in\mathbb{T}$, suppose that there exist two integers $a$ and $q$, such that
[…]
where $Q$ is a large positive integer. Under Hypothesis S,
$\sum_{p\leq x}\lambda_{\pi}(p)e(p^{k}\alpha)\log p\ll m^{2}\bigg{(}1+\frac{x^{k}}{qQ}\bigg{)}q^{\frac{1}{2}}x^{\delta+\varepsilon}(\log(x\mathfrak{q}_{\pi}))^{3}+x^{\frac{3}{4}}.$
Remark.
Comparing with previous similar results, Proposition 2.2 states the relationship between the upper bound of exponent sum of $\lambda_{\pi}(p)$ and $\pi$ for the first time. Hence, we can give some interesting applications.
Proposition 2.3.
Let $\pi\in\mathfrak{E}_{m}$, $\varepsilon$ be a sufficiently small positive number. For any real number $\alpha\in\mathbb{T}$, suppose that there exist two integers $a$ and $q$, such that
[…]
where $Q$ is a large positive integer. For all $A>0$ and $q\leq(\log x)^{A}$, under Hypothesis A,
$\sum_{p\leq x}(\lambda_{\pi\times\tilde{\pi}}(p)-1)e(p^{k}\alpha)\log p\ll_{\pi,A}\bigg{(}1+\frac{x^{k}}{qQ}\bigg{)}q^{\frac{m^{2}+1}{2}}x\exp\left(-c(\log x)^{1/4}\right),$
where $c$ is an ineffective constant depending on $A$ and $\pi$.
Note that the uniform bounds concerning exponential sums for all $\alpha\in\mathbb{T}$ are the key to obtaining (1.2) and (1.3). Although the nontrivial upper bound of exponential sums for higher ranks has not been settled uniformly, we apply a new approach to obtain not only (1.2) and (1.3), but also similar results for higher ranks. Consider Dirichlet rational approximation, i.e., there exist two integers $a$ and $q$ such that
[…]
hold for any $\alpha\in\mathbb{T}$, where $Q$ is a large positive integer. For $\alpha$ belonging to intervals around rational numbers $a/q$ with larger denominators, i.e., when $\alpha$ is in the so-called minor arcs, we can utilize some results of the following exponential sums (refer to Lemmas 5.15.3 below):
$\sum_{m\leq x}\Lambda(m)e\left(m^{k}\alpha\right),$
where $x$ is a large real number. In the other case, i.e., when $\alpha$ belongs to the so-called major arcs, we will use Propositions 2.12.3. By employing two distinct estimations, separately in major arcs and minor arcs, we can avoid the limitation of the previous methods concerning such problems to obtain Theorems 1.11.3.
3. Background on $L$-functions
3.1. Standard $L$-functions
Let $m\geq 1$ be an integer. Fix $\pi=\otimes_{p}\pi_{p}\in\mathfrak{E}_{m}$, where $p$ ranges over all primes. The standard function $L(s,\pi)$ is given by a Dirichlet series
$L(s,\pi)$ $=\prod_{p}L(s,\pi_{p})=\sum_{n=1}^{\infty}\frac{\lambda_{\pi}(n)}{n^{s}},$ (3.1)
with both the series and the product converging absolutely for $\mathrm{Re}(s)>1$. For each $p$, the local factor $L(s,\pi_{p})$ is defined in terms of Satake parameters $\{\alpha_{1,\pi}(p),...,\alpha_{n,\pi}(p)\}$ by
$L(s,\pi_{p}):=\prod_{j=1}^{m}\left(1-\frac{\alpha_{j,\pi}(p)}{p^{s}}\right)^{-1}=\sum_{i=0}^{\infty}\frac{\lambda_{\pi}\left(p^{i}\right)}{p^{is}}.$
Let $\mathfrak{q}_{\pi}$ denote the conductor of $\pi$. For all $j$, we have $\alpha_{j,\pi}(p)\neq 0$ when $p\nmid\mathfrak{q}_{\pi}$. When $p\mid\mathfrak{q}_{\pi}$, the case might be that $\alpha_{j,\pi}(p)=0$. At the archimedean place of $\mathbb{Q}$, there exist $m$ complex Langlands parameters $\mu_{\pi}(j)$ from which we define the Dirichlet series
$L_{\infty}(s,\pi):=\pi^{-ms/2}\prod_{j=1}^{m}\Gamma\left(\frac{s+\mu_{\pi}(j)}{2}\right).$ (3.2)
Let $\tilde{\pi}$ denote the contragredient representation of $\pi\in\mathfrak{E}_{m}$, which is also an irreducible cuspidal automorphic representation in $\mathfrak{E}_{m}$. For each $p\leq\infty$, we have the following equations
$\big{\{}\alpha_{j,\tilde{\pi}}(p):1\leq j\leq m\big{\}}=\big{\{}\overline{\alpha_{j,\pi}(p)}:1\leq j\leq m\big{\}}$
and
$\big{\{}\mu_{\tilde{\pi}}(j):1\leq j\leq m\big{\}}=\big{\{}\overline{\mu_{\pi}(j)}:1\leq j\leq m\big{\}}.$
For nontrivial $\pi$, the complete $L$-function is defined by
$\Lambda(s,\pi):=\mathfrak{q}_{\pi}^{\frac{s}{2}}L(s,\pi)L_{\infty}(s,\pi).$
The complete $L$-function $\Lambda(s,\pi)$ has an analytic continuation to the entire complex plane, and has finite order. Moreover, it satisfies a functional equation
$\Lambda(s,\pi)=W_{\pi}\Lambda(1-s,\tilde{\pi}),$
where $W_{\pi}$ is the root number (a complex number of magnitude 11 11). Due to Kim and Sarnak [23] $(2\leq m\leq 4)$ and Luo, Rudnick and Sarnak [26] $(m\geq 5)$, the best known record is
$\left|\alpha_{j,\pi}(p)\right|\leq p^{\theta_{m}}\quad\text{ and }\quad-\mathrm{Re}(\mu_{\pi}(j))\leq\theta_{m}$
for all primes $p$ and $1\leq j\leq m$, where
$\theta_{2}=\frac{7}{64},\quad\theta_{3}=\frac{5}{14},\quad\theta_{4}=\frac{9}{22},\quad\theta_{m}=\frac{1}{2}-\frac{1}{m^{2}+1}\ (m\geq 5).$ (3.3)
The generalized Ramanujan conjecture predicts that
$\theta_{m}=0$
for all $m\geq 1.$ The analytic conductor of $\pi$ is defined by
$C(\pi,t):=\mathfrak{q}_{\pi}\prod_{j=1}^{m}\left(1+\left|it+\mu_{\pi}(j)\right|\right),\quad C(\pi):=C(\pi,0).$ (3.4)
Taking the logarithmic derivative for $L(s,\pi)$, we obtain that, for $\mathrm{Re}(s)>1$,
$-\frac{L^{\prime}}{L}(s,\pi)=\sum_{p}\sum_{k=1}^{\infty}\frac{a_{\pi}(p^{k})\log p}{p^{ks}}=\sum_{n=1}^{\infty}\frac{a_{\pi}(n)\Lambda(n)}{n^{s}},$
where
$a_{\pi}\left(p^{k}\right)=\sum_{j=1}^{m}\alpha_{j,\pi}(p)^{k}.$ (3.5)
Lemma 3.1.
Let $\pi\in\mathfrak{E}_{m}$, $m\geq 2$, and $\varepsilon$ be a sufficiently small positive number.
  1. (1)
    Under Hypothesis S,
    $\sum_{n\leq x}\Lambda(n)a_{\pi}(n)\ll_{\varepsilon}mx^{\delta+\varepsilon}(\log x)^{2}\log\mathfrak{q}_{\pi}.$ (3.6)
  2. (2)
    Under Hypothesis S and GRC, the upper bound is improved to
    $\sum_{n\leq x}\Lambda(n)a_{\pi}(n)\ll x^{\delta}(\log x)^{2}(\log\log(x^{m}C(\pi)))^{2}.$ (3.7)
Proof.
By the Cauchy–Schwarz inequality, (3.15) and the Prime Number Theorem in short intervals [19], we can achieve a smoother expression by writing
$\displaystyle\sum_{n\leq x}\Lambda(n)a_{\pi}(n)=\sum_{n}\Lambda(n)a_{\pi}(n)\phi(n)+O(\sqrt{xy}),\quad x^{7/12}\ll y\ll x.$ (3.8)
Here, $\phi(z)$ is a function with support on the interval $[0,x+y]$, satisfying the conditions $\phi(z)=1$ for $1\leq z\leq x$ and $|\phi(z)|\leq 1$ elsewhere. The parameter $y$ will be determined later, subject to the inequality $1\leq y\leq x$. For example, we take
$\phi(z)=\min\left(\frac{z}{y},1,1+\frac{x-z}{y}\right)$
for $0\leq z\leq x+y$ and $\phi(z)=0$ if $z>x+y$. The Mellin transform of $\phi(z)$ satisfies
[…]
for $s=\sigma+it,\frac{1}{2}\leq\sigma\leq 2$. The sum in the right side of (3.16) can be expressed as
$\sum_{n}\Lambda(n)a_{\pi}(n)\phi(n)=\frac{1}{2\pi i}\int_{(2)}-\frac{L^{\prime}}{L}(s,\pi)\hat{\phi}(s)\mathrm{d}s,$
where the integral converges absolutely. Our operations will be conducted within the region defined as
$\mathcal{Z}:=\left\{s=\sigma+it\mid\sigma=\delta+\varepsilon\right\}.$
All the zeros of $L(s,\pi)$ are distanced from $\mathcal{Z}$ by at least $\varepsilon$. Therefore, it follows from [20, (5.27)] and [20, (5.28)] that for $s\in\mathcal{Z}$,
$-\frac{L^{\prime}}{L}(s,\pi)\ll\varepsilon^{-1}(\log C(\pi)+m\log(|t|+3)).$
By moving the integration from the line $\mathrm{Re}(s)=2$ to $\mathcal{Z}$ and combining the above inequality, we derive by Cauchy’s theorem that
$\displaystyle\sum_{n}\Lambda(n)a_{\pi}(n)\phi(n)$ $\displaystyle=\frac{1}{2\pi i}\int_{\mathcal{Z}}-\frac{L^{\prime}}{L}(s,\pi)\hat{\phi}(s)\mathrm{d}s$
$\displaystyle\ll\varepsilon^{-1}\int_{0}^{x/y}\frac{x^{\sigma}}{t+1}(\log C(\pi)+m\log(t+3))\mathrm{d}t$
$\displaystyle\quad+\varepsilon^{-1}\int_{x/y}^{\infty}\frac{x^{\sigma+1}}{t^{2}y}(\log C(\pi)+m\log(t+3))\mathrm{d}t$
$\displaystyle\ll\varepsilon^{-1}x^{\delta+\varepsilon}(m\log\mathfrak{q}_{\pi}+m\log(x/y+3))\log(x/y).$
We take $y=x\exp(-\sqrt{\log x})$ into the above inequality and combine (3.16), thereby obtain (3.6).
Under GRC, the difference is that we can use [20, (5.53)] to obtain that
$\sum_{n\leq x}\Lambda(n)a_{\pi}(n)=\sum_{|\gamma|\leq T}\frac{x^{\rho}-1}{\rho}+O\left(\frac{x}{T}(\log x)\log\left(x^{m}C(\pi)\right)\right),$
where $\rho=\beta+i\gamma$ runs over the zeros of $L(s,\pi)$ in the critical strip of height up to $T$ with $1\leq T\leq x$, and the implied constant is absolute. Using [20, (5.27)] with choosing $T=x^{1-\delta}(\log x)\log(x^{m}C(\pi))$, we find that
$\sum_{n\leq x}\Lambda(n)a_{\pi}(n)\ll x^{\delta}(\log x)^{2}(\log\log(x^{m}C(\pi)))^{2}.$
3.2. Rankin–Selberg $L$-functions
Let $\pi^{\prime}=\otimes_{p}\pi_{p}^{\prime}\in\mathfrak{E}_{m^{\prime}}$ $(m^{\prime}\geq 1)$ and $\pi=\otimes_{p}\pi_{p}\in\mathfrak{E}_{m}$ $(m\geq 1)$. The Rankin–Selberg $L$-function $L\left(s,\pi\times\pi^{\prime}\right)$ associated to $\pi$ and $\pi^{\prime}$ is of the form
$L\left(s,\pi\times\pi^{\prime}\right)=\prod_{p}L\left(s,\pi_{p}\times\pi_{p}^{\prime}\right)=\sum_{n=1}^{\infty}\frac{\lambda_{\pi\times\pi^{\prime}}(n)}{n^{s}}$
for $\mathrm{Re}(s)>1$. For each (finite) prime $p$, the inverse of the local factor $L\left(s,\pi_{p}\times\pi_{p}^{\prime}\right)$ is defined to be a polynomial in $p^{-s}$ of degree not exceeding $mm^{\prime}$,
$L\left(s,\pi_{p}\times\pi_{p}^{\prime}\right)^{-1}:=\prod_{j=1}^{m}\prod_{j^{\prime}=1}^{m^{\prime}}\left(1-\frac{\alpha_{j,j^{\prime},\pi\times\pi^{\prime}}(p)}{p^{s}}\right),$ (3.9)
where $\alpha_{j,j^{\prime},\pi\times\pi^{\prime}}(p)$ are suitable complex numbers. Given $\theta_{m}$ and $\theta_{m^{\prime}}$ as in (3.3), we have the pointwise bound
$\left|\alpha_{j,j^{\prime},\pi\times\pi^{\prime}}(p)\right|\leq p^{\theta_{m}+\theta_{m^{\prime}}}\leq p^{1-\frac{1}{mm^{\prime}}}.$ (3.10)
If $p\nmid\mathfrak{q}_{\pi}\mathfrak{q}_{\pi^{\prime}}$, we have the equality of sets
$\bigg{\{}\alpha_{j,j^{\prime},\pi\times\pi^{\prime}}(p):j\leq m,j^{\prime}\leq m^{\prime}\bigg{\}}=\bigg{\{}\alpha_{j,\pi}(p)\overline{\alpha_{j^{\prime},\pi^{\prime}}(p)}:j\leq m,j^{\prime}\leq m^{\prime}\bigg{\}}.$ (3.11)
At the archimedean place of $\mathbb{Q}$, there are $mm^{\prime}$ complex Langlands parameters $\mu_{\pi\times\pi^{\prime}}\left(j,j^{\prime}\right)$ from which we define the Dirichlet series
$L_{\infty}\left(s,\pi\times\pi^{\prime}\right):=\pi^{-\frac{mm^{\prime}s}{2}}\prod_{j=1}^{m}\prod_{j^{\prime}=1}^{m^{\prime}}\Gamma\left(\frac{s+\mu_{\pi\times\pi^{\prime}}(j,j^{\prime})}{2}\right).$
These parameters satisfy the pointwise bound
$\mathrm{Re}\left(\mu_{\pi\times\pi^{\prime}}\left(j,j^{\prime}\right)\right)\geq-\theta_{m}-\theta_{m^{\prime}}.$
Let $\mathfrak{q}_{\pi\times\pi^{\prime}}$ denote the conductor of $\pi\times\pi^{\prime}$. As with $L(s,\pi\times\pi^{\prime})$, we define the analytic conductor $C\left(\pi\times\pi^{\prime}\right)$ of $\pi\times\pi^{\prime}$ to be
$C\left(\pi\times\pi^{\prime},t\right):=\mathfrak{q}_{\pi\times\pi^{\prime}}\prod_{j=1}^{m}\prod_{j^{\prime}=1}^{m^{\prime}}\left(1+\left|it+\mu_{\pi\times\pi^{\prime}}\left(j,j^{\prime}\right)\right|\right),\ C\left(\pi\times\pi^{\prime}\right):=C\left(\pi\times\pi^{\prime},0\right).$ (3.12)
The combined work of Bushnell and Henniart [4] alongside [33, Appendix] establishes that
$C\left(\pi\times\pi^{\prime},t\right)\ll C\left(\pi\times\pi^{\prime}\right)(1+|t|)^{m^{\prime}m},\quad C\left(\pi\times\pi^{\prime}\right)\leq e^{O\left(m^{\prime}m\right)}C(\pi)^{m^{\prime}}C\left(\pi^{\prime}\right)^{m}.$ (3.13)
Let $r_{\pi\times\pi^{\prime}}\in\{0,1\}$ be the order of the pole of $L\left(s,\pi\times\pi^{\prime}\right)$ at $s=1$. We have $r_{\pi\times\pi^{\prime}}=1$ if and only if $\pi^{\prime}=\tilde{\pi}$. The function $\Lambda\left(s,\pi\times\pi^{\prime}\right)=(s(s-1))^{r_{\pi\times\pi^{\prime}}}\mathfrak{q}_{\pi\times\pi^{\prime}}^{s/2}L\left(s,\pi\right.$$\left.\times\pi^{\prime}\right)L_{\infty}\left(s,\pi\times\pi^{\prime}\right)$ is entire of order one. There exists a complex number $W\left(\pi\times\pi^{\prime}\right)$ of modulus one such that
$\Lambda\left(s,\pi\times\pi^{\prime}\right)=W\left(\pi\times\pi^{\prime}\right)\Lambda\left(1-s,\tilde{\pi}\times\tilde{\pi}^{\prime}\right).$
Furthermore, $L(s,\pi\times\tilde{\pi})$ extends to the complex plane with a simple pole at $s=1$, hence
$\sum_{n\leq x}\lambda_{\pi\times\widetilde{\pi}}(n)\sim x\mathrm{Res}_{s=1}L(s,\pi\times\widetilde{\pi})\ll x,$ (3.14)
as a result of a standard Tauberian argument. Taking the logarithmic derivative for $L(s,\pi\times\pi^{\prime})$, we obtain that, for $\mathrm{Re}(s)>1$,
$-\frac{L^{\prime}}{L}(s,\pi\times\pi^{\prime})=\sum_{p}\sum_{k=1}^{\infty}\frac{a_{\pi\times\pi^{\prime}}(p^{k})\log p}{p^{ks}}=\sum_{n=1}^{\infty}\frac{a_{\pi\times\pi^{\prime}}(n)\Lambda(n)}{n^{s}},$
where
$a_{\pi\times\pi^{\prime}}\left(p^{k}\right)=\sum_{j=1}^{m}\sum_{j^{\prime}=1}^{m^{\prime}}\alpha_{j,j^{\prime},\pi\times\pi^{\prime}}(p)^{k}.$
Building upon Shahidi’s non-vanishing result of $L(s,\pi\times\tilde{\pi})$ at $\mathrm{Re}(s)=1$ (see [32]), then
$\sum_{n\leq x}\Lambda(n)\left|a_{\pi}(n)\right|^{2}\ll\sum_{n\leq x}\Lambda(n)a_{\pi\times\widetilde{\pi}}(n)\sim x.$ (3.15)
Lemma 3.2.
Let $\pi\in\mathfrak{E}_{m}$. There exists a positive, absolute, and effectively computable constants $c_{1}$ such that the Rankin–Selberg L-function $L(s,\pi\times\widetilde{\pi})$ has at most one zero in the region
$\mathrm{Re}(s)\geq 1-\frac{c_{1}}{\log\left(C(\pi)^{m}(|\operatorname{Im}(s)|+3)^{m^{2}}\right)}.$
If such an exceptional zero $\beta_{1}$ exists, then it must be real and simple and there exists a positive, absolute, and effectively computable constants $c_{2}$ such that satisfy $\beta_{1}\leq 1-C(\pi)^{-c_{2}m}$.
Proof.
See [18, Theorem 2.1]. ∎
3.3. Twists
Let $\chi$ be a primitive Dirichlet character with conductor $q$. Fix $\pi=\otimes_{p}\pi_{p}\in\mathfrak{E}_{m}$ with $m\geq 2$. By [5], the twisted $L$-function $L(s,\pi\otimes\chi)$ equals
$\sum_{n=1}^{\infty}\frac{\lambda_{\pi\otimes\chi}(n)}{n^{s}}=\prod_{p}\prod_{j=1}^{m}\left(1-\frac{\alpha_{j,\pi\otimes\chi}(p)}{p^{s}}\right)^{-1}.$
Furthermore, by (3.11), if $p\nmid q$, then
$\left\{\alpha_{j,\pi\otimes\chi}(p):1\leq j\leq m\right\}=\left\{\chi(p)\alpha_{j,\pi}(p):1\leq j\leq m\right\}.$
Hence, for $\mathrm{Re}(s)>1$, we have
$\sum_{n=1}^{\infty}\frac{\chi(n)\lambda_{\pi}(n)}{n^{s}}$ $=\prod_{p}\prod_{j=1}^{m}\left(1-\frac{\chi(p)\alpha_{j,\pi}(p)}{p^{s}}\right)^{-1}=L(s,\pi\otimes\chi)\prod_{p\mid q}\prod_{j=1}^{m}\left(1-\frac{\alpha_{j,\pi\otimes\chi}(p)}{p^{s}}\right).$
In order to prove Proposition 2.12.3, we give some lemmas about zero-free regions and exceptional zeros.
Lemma 3.3.
Let $Q\geq 3$. There exists a constant $c_{\pi}>0$, depending effectively on $\pi\in\mathfrak{E}_{m}^{b}$, such that for all primitive Dirichlet characters $\chi(\bmod\thinspace q)$ with $q\leq Q$ with at most one exception, the $L$-function $L(s,\pi\otimes\chi)$ is non-zero in the region
$\mathrm{Re}(s)\geq 1-\frac{c_{\pi}}{\log(Q(3+|\mathrm{Im}(s)|))}.$
If the exceptional character $\chi_{1}$ exists, then $\chi_{1}$ is quadratic. $L\left(s,\pi\otimes\chi_{1}\right)$ has exactly one zero $\beta_{1}$ in this region, and $\beta_{1}$ is both real and simple. For all $\varepsilon_{0}>0$, there exists an ineffective constant $c_{\pi}(\varepsilon_{0})>0$ depending on $\pi$ and $\varepsilon_{0}$ such that $\beta_{1}\leq 1-$ $c_{\pi}(\varepsilon_{0})Q^{-\varepsilon_{0}}$.
Proof.
See [22, Theorem 4.1]. ∎
Lemma 3.4.
Let $\pi\in\mathfrak{E}_{m}$ and $\pi^{\prime}\in\mathfrak{E}_{m^{\prime}}$. For all $\varepsilon>0$, there exists an ineffective constant $c_{3}=c_{\pi,\pi^{\prime}}\left(\varepsilon\right)>0$ such that if $\chi\in\mathfrak{E}_{1},$
$\quad\left|L\left(\sigma+it,\pi\times\left(\pi^{\prime}\otimes\chi\right)\right)\right|\geq c_{3}C(it,\chi)^{-\varepsilon},\quad\sigma\geq 1-c_{3}C(it,\chi)^{-\varepsilon}$
where $C(it,\chi)\ll C(\chi)(|t|+1)$, $C(\chi)$ is defined in (3.4).
Proof.
See [11, Theorem 1.1]. ∎
Lemma 3.5.
Let $\pi\in\mathfrak{E}_{m}$, $m\geq 2$, $\chi$ be a nontrivial primitive Dirichlet character, and $A$ be an arbitrary positive constant. Under Hypothesis A, there exists an ineffective constant $c_{4}>0$ such that
$\sum_{n\leq x}\Lambda(n)a_{\pi\times\tilde{\pi}\otimes\chi}(n)\ll_{\pi}x\exp(-c_{4}C(\chi)^{-1/2A}\sqrt{\log x}).$
Proof.
By the Cauchy–Schwarz inequality, Hypothesis A and the Prime Number Theorem in short intervals [19], we can achieve a smoother expression by writing
$\displaystyle\sum_{n\leq x}\Lambda(n)a_{\pi\times\tilde{\pi}\otimes\chi}(n)=\sum_{n}\Lambda(n)a_{\pi\times\tilde{\pi}\otimes\chi}(n)\phi(n)+O(\sqrt{xy}),\quad x^{7/12}\ll y\ll x.$ (3.16)
Here, $\phi(z)$ is the same function in Lemma 3.1. By the same proof of Lemma 3.1, the sum in the right side of (3.16) can be expressed as
$\sum_{n}\Lambda(n)a_{\pi\times\tilde{\pi}\otimes\chi}(n)\phi(n)=\frac{1}{2\pi i}\int_{(2)}-\frac{L^{\prime}}{L}(s,\pi\times\tilde{\pi}\otimes\chi)\hat{\phi}(s)\mathrm{d}s,$
where the integral converges absolutely. Our operations will be conducted within the region defined as
$\mathcal{Z}:=\left\{s=\sigma+it\mid\sigma=1-c_{3}C(it,\chi)^{-\varepsilon}/2\right\},$
where $c_{3}$ is defined in the above lemma. All the zeros of $L(s,\pi)$ are distanced from $\mathcal{Z}$ by at least $c_{3}C(it,\chi)^{-\varepsilon}/2$. Therefore, it follows from [20, (5.27)] and [20, (5.28)] that for $s\in\mathcal{Z}$,
$-\frac{L^{\prime}}{L}(s,\pi\times\tilde{\pi}\otimes\chi)\ll\frac{c_{3}}{2}C(it,\chi)^{-\varepsilon}(\log C(\pi)+m\log(|t|+3)).$
By moving the integration from the line $\mathrm{Re}(s)=2$ to $\mathcal{Z}$ and combining the above inequality, we derive by Cauchy’s theorem that
$\displaystyle\sum_{n}\Lambda(n)a_{\pi\times\tilde{\pi}\otimes\chi}(n)\phi(n)$ $\displaystyle=\frac{1}{2\pi i}\int_{\mathcal{Z}}-\frac{L^{\prime}}{L}(s,\pi\times\tilde{\pi}\otimes\chi)\hat{\phi}(s)\mathrm{d}s$
$\displaystyle\ll\int_{0}^{x/y}\frac{x^{\sigma}}{t+1}(\frac{c_{3}}{2}(C(\chi)(|t|+1))^{-\varepsilon}(\log C(\pi)+m\log(|t|+3)))\mathrm{d}t$
$\displaystyle\quad+\int_{x/y}^{\infty}\frac{x^{\sigma+1}}{t^{2}y}(\frac{c_{3}}{2}(C(\chi)(|t|+1))^{-\varepsilon}(\log C(\pi)+m\log(|t|+3)))\mathrm{d}t.$
We take $y=x/(\log x)^{A}$ and $\varepsilon=1/(2A)$ into the above inequality and combine (3.16), thereby
$\displaystyle\sum_{n}\Lambda(n)a_{\pi\times\tilde{\pi}\otimes\chi}(n)\phi(n)\ll_{\pi}x\exp(-c_{4}C(\chi)^{-1/2A}\sqrt{\log x}),$
where $c_{4}=c_{3}/4$ and the implied constant is dependent on $\pi$. ∎
4. Proof of Propositions 2.12.3
4.1. Reduction
Define the sum
$\sum_{n\leq x}\Lambda(n)g(n)e(n^{k}\alpha),$
where $g(n)$ is an arithmetic function. According to the Dirichlet approximation theorem, there exist integers $a$ and $q$ that make any $\alpha\in[0,1)$ be expressed as (2.4), where $Q$ is a parameter to be chosen later. Combining (2.4) with partial summation, the above sum is bounded by
$\left(1+\frac{x^{k}}{qQ}\right)\max_{1\leq t\leq x}\left|\sum_{n\leq t}\Lambda(n)g(n)e\left(\frac{an^{k}}{q}\right)\right|.$ (4.1)
Due to the orthogonality of characters, we find that
$\sum_{n\leq t}\Lambda(n)g(n)e\left(\frac{an^{k}}{q}\right)$ $=\sum_{d\mid q}\sum_{\begin{subarray}{c}h=1\\ (h,q/d)=1\end{subarray}}^{q}e\left(\frac{ad^{k}h^{k}}{q}\right)\sum_{\begin{subarray}{c}l\leq t/d\\ l\equiv h(\bmod q/d)\end{subarray}}\Lambda(dl)g(dl)$ (4.2)
$=\sum_{d\mid q}\sum_{h=1}^{q/d}e\left(\frac{ad^{k}h^{k}}{q}\right)\sum_{l\leq t/d}\Lambda(dl)g(dl)\frac{1}{\varphi(q/d)}\sum_{\begin{subarray}{c}\chi(\bmod q/d)\\ (h,q/d)=1\end{subarray}}\bar{\chi}(h)\chi(l)$
$\ll\sum_{d\mid q}\frac{1}{\varphi(q/d)}\max_{\chi\left(\bmod q/d\right)}\bigg{|}\sum_{l\leq t/d}\Lambda(dl)g(dl)\chi(l)\bigg{|}$
×$\quad\times\sum_{\chi\left(\bmod q/d\right)}\bigg{|}\sum_{\begin{subarray}{c}h=1\\ (h,q/d)=1\end{subarray}}^{q/d}\bar{\chi}(h)e\left(\frac{ad^{k}h^{k}}{q}\right)\bigg{|},$
where $\varphi$ is the Euler function. By the Cauchy–Schwarz inequality and Parseval’s identity, the innermost sum over $\chi$ is bounded by
$\displaystyle\varphi\left(q/d\right)^{1/2}\left(\sum_{\chi\left(\bmod q/d\right)}\bigg{|}\sum_{\begin{subarray}{c}h=1\\ (h,q/d)=1\end{subarray}}^{q/d}\bar{\chi}(h)e\left(\frac{ad^{k}h^{k}}{q}\right)\bigg{|}^{2}\right)^{1/2}$
$\displaystyle$== $\displaystyle\varphi\left(q/d\right)^{1/2}\left(\sum_{\begin{subarray}{c}h_{1}=1\\ \left(h_{1},q/d\right)=1\end{subarray}}^{q/d}\sum_{\begin{subarray}{c}h_{2}=1\\ \left(h_{2},q/d\right)=1\end{subarray}}^{q/d}e\left(\frac{ad^{k}\left(h_{1}^{k}-h_{2}^{k}\right)}{q}\right)\sum_{\begin{subarray}{c}\chi\left(\bmod q/d\right)\end{subarray}}\bar{\chi}\left(h_{1}\right)\chi\left(h_{2}\right)\right)^{1/2}$
$\displaystyle$== $\displaystyle\varphi\left(q/d\right)\left(\sum_{\begin{subarray}{c}h_{1}=1\\ \left(h_{1},q/d\right)=1\end{subarray}}^{q/d}\sum_{\begin{subarray}{c}h_{2}=1\\ h_{2}\equiv h_{1}\left(\bmod q/d\right)\end{subarray}}^{q/d}e\left(\frac{ad^{k}\left(h_{1}^{k}-h_{2}^{k}\right)}{q}\right)\right)^{1/2}=\varphi\left(q/d\right)^{3/2}.$ (4.3)
Inserting the estimates (4.2) and (4.1) into (4.1), then
$\sum_{n\leq x}\Lambda(n)g(n)e\left(n^{k}\alpha\right)\ll\left(1+\frac{x^{k}}{qQ}\right)\max_{1\leq t\leq x}\left\{q^{1/2}\sum_{d\mid q}\frac{1}{d^{1/2}}\max_{\chi(\bmod q/d)}\bigg{|}\sum_{l\leq t/d}\Lambda(dl)g(dl)\chi(l)\bigg{|}\right\}.$ (4.4)
4.2. Proof of Proposition 2.1
It follows from (3.1) and (3.5) that
$\sum_{p\leq x}\lambda_{\pi}(p)e(p^{k}\alpha)\log p=\sum_{n\leq x}\Lambda(n)a_{\pi}(n)e(n^{k}\alpha)-\sum_{p^{k_{1}}\leq x,k_{1}\geq 2}a_{\pi}(p^{k_{1}})e(p^{k_{1}}\alpha)\log p.$ (4.5)
Using the Cauchy–Schwarz inequality, (3.14) and (3.15), the left side of the above equation is equal to
$\sum_{n\leq x}\Lambda(n)a_{\pi}(n)e(n^{k}\alpha)+O(x^{3/4}).$ (4.6)
Let $g(n)=a_{\pi}(n)$ in (4.4), the above summation is
$\ll\left(1+\frac{x^{k}}{qQ}\right)\max_{1\leq t\leq x}\left\{q^{1/2}\sum_{d\mid q}\frac{1}{d^{1/2}}\max_{\chi(\bmod q/d)}\bigg{|}\sum_{l\leq t/d}\Lambda(dl)a_{\pi}(dl)\chi(l)\bigg{|}\right\}.$ (4.7)
Let $l=l_{1}l_{2}$ with $l_{1}\mid d^{\infty}$ and $\left(l_{2},l_{1}d\right)=1.$ By the multiplicative property of $a_{\pi}(dl)$, $l_{1}\mid d^{\infty}$ and $\left(l_{2},l_{1}d\right)=1$, the formula in the brace is
[…]
$\displaystyle+q^{1/2}\max_{\chi(\bmod q)}\bigg{|}\sum_{l_{2}\leq t}\Lambda(l_{2})a_{\pi}\left(l_{2}\right)\chi\left(l_{2}\right)\bigg{|}.$ (4.8)
Due to the property of $\Lambda(n)$ and the upper bound of $a_{\pi}(n)$, the first item on the right side of the above formula is
$\displaystyle\ll q^{1/2}t^{1/2}\sum_{p^{k}\mid q,\ k\geq 1}\frac{1}{p^{k/2}}\bigg{|}\sum_{p^{k_{1}}\leq t/p^{k},\ k_{1}\geq 1}\log p\bigg{|}\ll(qt)^{1/2}(\log q)\log t.$ (4.9)
For the second item on the right side of (4.2), if $\chi^{\prime}(\bmod\thinspace q^{\prime})$ is the primitive Dirichlet character that induces $\chi(\bmod\thinspace q)$, then $q^{\prime}|q$. By (3.3), we get
$\left|\sum_{l_{2}\leq t}\Lambda(l_{2})a_{\pi}\left(l_{2}\right)\chi\left(l_{2}\right)-\sum_{l_{2}\leq t}\Lambda(l_{2})a_{\pi\otimes\chi^{\prime}}(l_{2})\right|\leq\sum_{p\mid q}\sum_{p^{k}\leq t,\ k\geq 1}\left|a_{\pi}\left(p^{k}\right)\right|\log p\ll_{\varepsilon}t^{\theta_{m}+\varepsilon}.$ (4.10)
Without loss of generality, assume that $\chi^{\prime}$ is the exceptional character $\chi_{1}$ in Lemma 3.3, and let $\beta^{\prime}$ be the corresponding exceptional zero. It follows from (3.13) and (3.12) that the modulus of the conductor of $\pi\otimes\chi^{\prime}$ is smaller than $\mathfrak{q}_{\pi}q^{\prime m}$, we apply [20, (5.52)] to the twisted $L$-function $L(s,\pi\otimes\chi^{\prime})$ and conclude that
$\sum_{n\leq x}\Lambda(n)a_{\pi\otimes\chi^{\prime}}(n)=-\frac{x^{\beta^{\prime}}}{\beta^{\prime}}+O\bigg{(}\mathfrak{q}_{\pi}^{\frac{1}{2}}q^{\prime\frac{m}{2}}x\exp\left(-\frac{c_{\pi}}{2}\sqrt{\log x}\right)\bigg{)}.$
Since Lemma 3.3 gives the bound $\beta^{\prime}\leq 1-c_{\pi}(\varepsilon_{0})q^{-\varepsilon_{0}}$ where $\varepsilon_{0}>0$ is a fixed constant chosen later, for all $q\leq(\log t)^{A}$, there holds
$\displaystyle\sum_{l_{2}\leq t}\Lambda(l_{2})a_{\pi\otimes\chi^{\prime}}(l_{2})\ll_{\pi}t\exp\left(-c_{\pi}(\varepsilon_{0})(\log t)q^{-\varepsilon_{0}}\right)+tq^{\frac{m}{2}}\exp\left(-\frac{c_{\pi}}{4}\sqrt{\log t}\right).$
Choose $\varepsilon_{0}=1/(2A)$. Then inserting (4.7)–(4.10) and the above inequality into (4.6), we get
$\sum_{p\leq x}\lambda_{\pi}(p)e(p^{k}\alpha)\log p\ll_{\pi,A}\bigg{(}1+\frac{x^{k}}{qQ}\bigg{)}q^{\frac{m+1}{2}}x\exp\left(-c\sqrt{\log x}\right),$
where $c=\min\left\{\frac{c_{\pi}}{8},\frac{c_{\pi}(1/(2A))}{2}\right\}$.
4.3. Proof of Proposition 2.2
Lemma 3.1 gives that
$\sum_{l_{2}\leq t}\Lambda(l_{2})a_{\pi\otimes\chi^{\prime}}(l_{2})\ll mt^{\delta+\varepsilon}(\log t)^{2}\log(\mathfrak{q}_{\pi}q^{m}).$
Inserting (4.7)–(4.10) and the above inequality into (4.6), we get
$\sum_{p\leq x}\lambda_{\pi}(p)e(p^{k}\alpha)\log p\ll m^{2}\bigg{(}1+\frac{x^{k}}{qQ}\bigg{)}q^{\frac{1}{2}}x^{\delta+\varepsilon}(\log x\mathfrak{q}_{\pi})^{3}+x^{3/4}.$
4.4. Proof of Proposition 2.3
Under Hypothesis A, we utilize the Cauchy–Schwarz inequality to obtain
$\sum_{p\leq x}(\lambda_{\pi\times\tilde{\pi}}(p)-1)e(p^{k}\alpha)\log p=\sum_{n\leq x}\Lambda(n)(a_{\pi\times\tilde{\pi}}(n)-1)e(n^{k}\alpha)+O(x^{3/4}).$ (4.11)
Let $g(n)=a_{\pi\times\tilde{\pi}}(n)-1$ in (4.4). Then we get
$\displaystyle\sum_{n\leq x}$ $\displaystyle\Lambda(n)(a_{\pi\times\tilde{\pi}}(n)-1)e(n^{k}\alpha)$
[…]
Use the relationship between $\tilde{\pi}\otimes\chi$ and $\tilde{\pi}$, we divide the summation in the brace and obtain that it is
[…]
We use [20, (5.52)] and Lemma 3.2 to get
$\displaystyle\sum_{l_{2}\leq t}\Lambda(l_{2})a_{\pi\times\tilde{\pi}}\left(l_{2}\right)$ $\displaystyle=t+O\left(t\exp(-C(\pi)^{-c_{2}m}\log t)+t\mathfrak{q}_{\pi\times\tilde{\pi}}^{1/2}\exp\left(-\frac{c_{1}}{2m^{2}}\sqrt{\log t}\right)\right),$
where $c_{2}$ is a computable constant in Lemma 3.2 and $c_{1}$ in Lemma 3.2. Then we obtain that
$\displaystyle\sum_{l_{2}\leq t}\Lambda(l_{2})a_{\pi\times\tilde{\pi}}\left(l_{2}\right)=t+tq^{\frac{m^{2}}{2}}\exp\left(-c_{5}\sqrt{\log t}\right),$ (4.14)
where $c_{5}$ is a computable constant depending on $c_{2}$, $c_{1}$ and $\pi$. By the zero-free region of Vinogradov-Korobov [20, Corollary 8.30], we obtain
$\sum_{l_{2}\leq t}\Lambda(l_{2})=t+O\left\{t\exp\left(-c_{6}\frac{(\log t)^{3/5}}{(\log\log t)^{1/5}}\right)\right\},$ (4.15)
where $c_{6}$ is a positive number. For $\chi$ satisfying $\tilde{\pi}\otimes\chi\neq\tilde{\pi}$, there exists $\chi^{\prime}(\bmod\ q^{\prime})$ is the primitive Dirichlet character that induces $\chi(\bmod\thinspace q)$, and $q^{\prime}|q$. By Lemma 3.5, for all $q\leq(\log t)^{A/2}$, there holds
$\displaystyle\sum_{l_{2}\leq t}\Lambda(l_{2})a_{\pi\times\tilde{\pi}\otimes\chi^{\prime}}\left(l_{2}\right)\ll t\exp(-c_{4}C(\chi)^{-1/(2A)}\sqrt{\log t})\ll t\exp(-c_{7}(\log t)^{1/4}),$
where $c_{7}=c_{4}/2$ and $c_{4}$ is an ineffective constant in Lemma 3.5. Due to the Siegel–Walfisz theorem, for all $q\leq(\log t)^{A/2}$, there exists a positive absolute constant $c_{8}$ such that
$\displaystyle\sum_{l_{2}\leq t}\Lambda(l_{2})\chi^{\prime}(l_{2})\ll t\exp\left(-c_{8}\frac{(\log t)^{3/5}}{(\log\log t)^{1/5}}\right).$ (4.16)
Inserting (4.4)–(4.16) into (4.11), we get
$\sum_{p\leq x}(\lambda_{\pi\times\tilde{\pi}}(p)-1)e(p^{k}\alpha)\log p\ll_{\pi,A}\bigg{(}1+\frac{x^{k}}{qQ}\bigg{)}q^{\frac{m^{2}+1}{2}}x\exp\left(-c(\log x)^{1/4}\right),$
where $c=c_{7}$.
5. Proof of Theorems 1.11.3
5.1. Preliminaries and lemmas
The initial step of the Hardy–Littlewood circle method would be to divide the unit interval into the major arcs and the minor arcs. The major arcs $\mathfrak{M}$ and minor arcs $\mathfrak{m}$ are defined as follows:
$\ \ \ \ \ \ \ \ \mathfrak{M}=\bigcup_{q\leq P}\bigcup_{1\leq a\leq q\atop(a,q)=1}\mathfrak{M}(q,a),\ \ \ \mathfrak{M}(q,a)=\Big{[}\frac{a}{q}-\frac{1}{qQ},\frac{a}{q}+\frac{1}{qQ}\Big{]},\ \ \ \ \ \mathfrak{m}=\big{[}1/Q,1+1/Q\big{]}\setminus\mathfrak{M},$ (5.1)
where $1<2P<Q$, $PQ=N$, $P$, $Q$ are parameters to be chosen later. Let
$\displaystyle T_{k,1}(N,\alpha)=\sum_{p_{1}\leq N^{1/k}}\lambda_{\pi}(p_{1})e(p_{1}^{k}\alpha),\quad H_{k,u^{\prime}}(N,\alpha)=\sum_{p_{u^{\prime}}\leq N^{1/k}}e(p_{u^{\prime}}^{k}\alpha)\quad(2\leq u^{\prime}\leq u).$ (5.2)
Using Parseval’s identity, we write
$\displaystyle\sum_{(p_{1},p_{2},\dots,p_{u})\in\mathcal{J}_{k,u}(N)}\lambda_{\pi}(p_{1})$ $\displaystyle=\int_{0}^{1}T_{k,1}(N,\alpha)\prod_{u^{\prime}=2}^{u}H_{k,u^{\prime}}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
$\displaystyle=\int_{\mathfrak{M}}T_{k,1}(N,\alpha)\prod_{u^{\prime}=2}^{u}H_{k,u^{\prime}}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
$\displaystyle\quad+\int_{\mathfrak{m}}T_{k,1}(N,\alpha)\prod_{u^{\prime}=2}^{u}H_{k,u^{\prime}}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha.$ (5.3)
To control the integral on the minor arcs, we need certain estimates concerning exponential sums over primes.
Lemma 5.1.
Suppose that $\alpha$ is a real number and there exist integers $a$ and $q$ such that $\alpha$ satisfies
$\alpha=\frac{a}{q}+\lambda,\quad 1\leq a\leq q,\quad(a,q)=1,\quad|\lambda|\leq\frac{1}{q^{2}}.$ (5.4)
  1. (1)
    For $k=1$,
    $\sum_{m\leq x}\Lambda(m)e(m^{k}\alpha)\ll x(\log x)^{3}\left(\frac{1}{q}+\frac{1}{x^{2/5}}+\frac{q}{x}\right)^{1/2}.$
  2. (2)
    For all $k\geq 2$ and $\varepsilon_{1}>0$,
    $\sum_{m\leq x}\Lambda(m)e(m^{k}\alpha)\ll x^{1+\varepsilon_{1}}\left(\frac{1}{q}+\frac{1}{x^{1/2}}+\frac{q}{x^{k}}\right)^{1/K^{2}},$
    where $K=2^{k-1}$.
Proof.
For $k=1$ one can see [38], for $k\geq 2$ one can see [13, Theorem 1]. ∎
Sometimes, for $k\geq 2$ it is better to have the factor $x^{\varepsilon_{1}}$ above replaced by some power of $\log x$. This can be done by a simple modification of the above argument.
Lemma 5.2.
Let $k\geq 2$. Suppose that $\alpha$ is a real number and there exist integers $a$ and $q$ such that $\alpha$ satisfies (5.4). Then
$\sum_{m\leq x}\Lambda(m)e(m^{k}\alpha)\ll x(\log x)^{c^{\prime}}\left(\frac{1}{q}+\frac{1}{x^{1/2}}+\frac{q}{x^{k}}\right)^{1/(2K^{2})},$
where $c^{\prime}$ is a positive constant that depends only on $k$, $K$ is defined in Lemma 5.1.
Lemma 5.3.
Let $k\geq 1$, $\gamma_{k}=1/2+\log k/\log 2$. Suppose that $\alpha$ is a real number and there exist integers $a$ and $q$ such that $\alpha$ satisfies (5.4). Then
$\sum_{m\leq x}\Lambda(m)e(m^{k}\alpha)\ll(d(q))^{\gamma_{k}}\left\{x^{1/2}\sqrt{q\left(1+|\lambda|x^{k}\right)}+x^{4/5}+\frac{x}{\sqrt{q\left(1+|\lambda|x^{k}\right)}}\right\}(\log x)^{c^{\prime\prime}},$
where $d(n)$ is the divisor function and $c^{\prime\prime}$ is an absolute positive constant.
Proof.
See [31, Theorem 1.1]. ∎
5.2. Proof of Theorem 1.1
5.2.1. Major arcs
As for $r_{\pi}(N)$, we use the Polya–Vinogradov inequality to derive that
$\int_{0}^{1}\left|H_{1,u^{\prime}}(N,\alpha)\right|^{2}\mathrm{d}\alpha\leq\sum_{p_{u^{\prime}_{1}}\leq N}\sum_{p_{u^{\prime}_{2}}\leq N}\int_{0}^{1}e((p_{u^{\prime}_{1}}-p_{u^{\prime}_{2}})\alpha)\mathrm{d}\alpha\leq\sum_{p\leq N}1\ll N/\log N.$ (5.5)
Therefore, by applying the Cauchy–Schwarz inequality, the above inequality, partial summation and Proposition 2.1,
$\displaystyle\int_{\mathfrak{M}}$ $\displaystyle T_{1,1}(N,\alpha)H_{1,2}(N,\alpha)H_{1,3}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
$\displaystyle\ll\left(\max_{\alpha\in\mathfrak{M}}\left|T_{1,1}(N,\alpha)\right|\right)\left\{\int_{0}^{1}\left|H_{1,2}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}\left\{\int_{0}^{1}\left|H_{1,3}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}$
$\displaystyle\ll_{\pi,B}P^{\frac{m+1}{2}}N^{2}\exp\left(-c\sqrt{\log N}\right),$ (5.6)
where $c$ is the constant in Proposition 2.1 and $P\leq(\log N)^{100B}$.
Regarding $w_{\pi}(N)$, we apply [3, Theorem 3] to give
$\displaystyle\int_{0}^{1}\left|H_{2,2}(N,\alpha)H_{2,3}(N,\alpha)\right|^{2}\mathrm{d}\alpha\leq\sum_{p_{2}^{2}+p_{3}^{2}=p_{2}^{\prime 2}+p_{3}^{\prime 2}\atop p_{2},p_{3},p_{2}^{\prime},p_{3}^{\prime}\leq N^{1/2}}1\ll N\log N.$ (5.7)
Hence it follows from the Cauchy–Schwarz inequality, the above inequality, partial summation and Proposition 2.1 that
$\displaystyle\int_{\mathfrak{M}}T_{2,1}(N,\alpha)$ $\displaystyle H_{2,2}(N,\alpha)H_{2,3}(N,\alpha)H_{2,4}(N,\alpha)H_{2,5}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
[…]
×$\displaystyle\quad\times\left\{\int_{0}^{1}\left|H_{2,4}(N,\alpha)H_{2,5}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}$
[…]
5.2.2. Minor arcs
By the Polya–Vinogradov inequality and [20, (5.48)],
$\int_{0}^{1}\left|T_{k,1}(N,\alpha)\right|^{2}\mathrm{d}\alpha$ $\leq\sum_{p_{1}\leq N^{1/k}}\sum_{p_{2}\leq N^{1/k}}\lambda_{\pi}(p_{1})\bar{\lambda}_{\pi}(p_{2})\int_{0}^{1}e((p_{1}^{k}-p_{2}^{k})\alpha)\mathrm{d}\alpha$ (5.9)
$\leq\sum_{p\leq N^{1/k}}|\lambda_{\pi}(p)|^{2}\ll N^{1/k}m^{2}(\log(N\mathfrak{q}_{\pi}))^{2}.$
Following from the Cauchy–Schwarz inequality, (5.5) and the above inequality, we have
$\displaystyle\int_{\mathfrak{m}}$ $\displaystyle T_{1,1}(N,\alpha)H_{1,2}(N,\alpha)H_{1,3}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
$\displaystyle\leq\left(\max_{\alpha\in\mathfrak{m}}\left|H_{1,2}(N,\alpha)\right|\right)\left\{\int_{0}^{1}\left|T_{1,1}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}\left\{\int_{0}^{1}\left|H_{1,3}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}$
$\displaystyle\ll\left(\max_{\alpha\in\mathfrak{m}}\left|H_{1,2}(N,\alpha)\right|\right)Nm\log(N\mathfrak{q}_{\pi}).$ (5.10)
Using the first inequality in Lemma 5.1 and partial summation, $H_{1,2}(N,\alpha)$ over the minor arcs is bounded by
$(Q^{\frac{1}{2}}N^{\frac{1}{2}}+P^{-\frac{1}{2}}N+N^{\frac{4}{5}})(\log N)^{3}.$
Inserting this bound into (5.2.2), then
$\int_{\mathfrak{m}}$ $T_{1,1}(N,\alpha)H_{1,2}(N,\alpha)H_{1,3}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha\ll N^{2}m\log(N\mathfrak{q}_{\pi})(P^{-\frac{1}{2}}+N^{-\frac{1}{5}})(\log N)^{3}.$ (5.11)
For $w_{\pi}(N)$, we obtain that
$\displaystyle\int_{0}^{1}\left|T_{2,1}(N,\alpha)H_{2,3}(N,\alpha)\right|^{2}\mathrm{d}\alpha$ $\displaystyle\ll\sum_{p_{1}^{2}+p_{3}^{2}=p_{1}^{\prime 2}+p_{3}^{\prime 2}\atop p_{1},p_{3},p_{1}^{\prime},p_{3}^{\prime}\leq N^{\frac{1}{2}}}|\lambda_{\pi}(p_{1})\lambda_{\pi}(p_{1}^{\prime})|$
$\displaystyle\ll\sum_{p_{1}\leq N^{\frac{1}{2}}}|\lambda_{\pi}(p_{1})|^{2}\sum_{p_{1}^{\prime}\leq N^{\frac{1}{2}}}\sum_{p_{1}^{2}-p_{1}^{\prime 2}=(p_{3}^{\prime}+p_{3})(p_{3}^{\prime}-p_{3})\atop p_{3},p_{3}^{\prime}\leq N^{\frac{1}{2}}}1$
$\displaystyle\ll Nm^{2}\log^{2}(N\mathfrak{q}_{\pi})\log N.$ (5.12)
Treat $w_{\pi}(N)$ similarly as the estimate of $r_{\pi}(N)$ on the minor arcs to get
$\displaystyle\int_{\mathfrak{m}}T_{2,1}(N,\alpha)H_{2,2}(N,\alpha)H_{2,3}(N,\alpha)H_{2,4}(N,\alpha)H_{2,5}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
$\displaystyle\leq\left(\max_{\alpha\in\mathfrak{m}}\left|H_{2,2}(N,\alpha)\right|\right)\left\{\int_{0}^{1}\left|T_{2,1}(N,\alpha)H_{2,3}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}$
$\displaystyle\quad\cdot\left\{\int_{0}^{1}\left|H_{2,4}(N,\alpha)H_{2,5}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}$
$\displaystyle\ll\left(\max_{\alpha\in\mathfrak{m}}\left|H_{2,2}(N,\alpha)\right|\right)Nm\log(N\mathfrak{q}_{\pi})\log N.$ (5.13)
In accordance with Lemma 5.2 and by partial summation, we give the following upper bound
$\max_{\alpha\in\mathfrak{m}}\left|H_{2,2}(N,\alpha)\right|\ll N^{1/2}(\log N)^{c^{\prime}}\left(\frac{1}{P}+\frac{1}{N^{1/4}}+\frac{Q}{N}\right)^{1/8}.$
Inserting the above estimate into (5.2.2),
$\displaystyle\int_{\mathfrak{m}}T_{2,1}(N,\alpha)H_{2,2}(N,\alpha)H_{2,3}(N,\alpha)H_{2,4}(N,\alpha)H_{2,5}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
$\displaystyle\ll N^{3/2}m\log(N\mathfrak{q}_{\pi})(\log N)^{c^{\prime}+1}\bigg{(}P^{-1/8}+N^{-1/32}\bigg{)}.$ (5.14)
5.2.3. Choosing prameters
For $r_{\pi}(N)$, let $P=(\log N)^{2B+8}$. Then by (5.2.1), (5.11) and (5.1), we have
$r_{\pi}(N)=\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\lambda_{\pi}\left(p_{1}\right)\ll_{\pi,B}N^{2}(\log N)^{-B}.$
For $w_{\pi}(N)$, we choose $P=(\log N)^{8B+8c^{\prime}+16}$. It follows from (5.2.1), (5.2.2) and (5.1) that
$w_{\pi}(N)=\sum_{(p_{1},p_{2},p_{3},p_{4},p_{5})\in\mathcal{J}_{2,5}(N)}\lambda_{\pi}\left(p_{1}\right)\ll_{\pi,B}N^{3/2}(\log N)^{-B}.$
5.3. Proof of Theorem 1.2
We use Proposition 2.2 to improve the upper bounds in (5.2.1) and (5.2.1) respectively. Then
$\int_{\mathfrak{M}}$ $T_{1,1}(N,\alpha)H_{1,2}(N,\alpha)H_{1,3}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha\ll Pm^{2}N^{1+\delta+\varepsilon}(\log(N\mathfrak{q}_{\pi}))^{3}+N^{7/4}/\log N$ (5.15)
and
$\displaystyle\int_{\mathfrak{M}}$ $\displaystyle T_{2,1}(N,\alpha)H_{2,2}(N,\alpha)H_{2,3}(N,\alpha)H_{2,4}(N,\alpha)H_{2,5}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
$\displaystyle\ll Pm^{2}N^{1+\delta/2+\varepsilon}(\log(N\mathfrak{q}_{\pi}))^{3}\log N+N^{11/8}\log N.$ (5.16)
For $r_{\pi}(N)$ under Hypothesis S, we choose $P=N^{\frac{2}{3}(1-\delta)}$ in (5.11) and (5.15) to get
$r_{\pi}(N)\ll m^{2}N^{5/3+\delta/3+\varepsilon}(\log(N\mathfrak{q}_{\pi}))^{3}.$
For $w_{\pi}(N)$ under Hypothesis S, put $P=N^{\frac{1}{3}(1-\delta)},\ Q=N/P$ and $P^{\prime}=N^{1/3},\ Q^{\prime}=N/P^{\prime}$. On the minor arcs, one notes that each $\alpha\in\mathfrak{m}$ can be written as (5.1) with $P<q\leq Q$. By Dirichlet’s rational approximation, each $\alpha\in\mathfrak{m}$ can also be written as (2.4) with $1\leq q\leq Q^{\prime}$ and $|\lambda|\leq 1/qQ^{\prime}$. Let $\mathfrak{n}$ be the subset of $\alpha\in\mathfrak{m}$ such that $P^{\prime}<q\leq Q^{\prime}$ and $|\lambda|\leq 1/qQ^{\prime}$. We now apply the second inequality of Lemma 5.1 and partial summation to get
$\max_{\alpha\in\mathfrak{n}}\left|H_{2,2}(N,\alpha)\right|\ll N^{1/2+\varepsilon}\left(\frac{1}{P^{\prime}}+\frac{1}{N^{1/4}}+\frac{Q^{\prime}}{N}\right)^{1/4}\ll N^{7/16+\varepsilon}.$ (5.17)
For $\alpha\in\mathfrak{m}\backslash\mathfrak{n}$, we have
$P<q\leq P^{\prime},\quad|\lambda|\leq\frac{1}{qQ^{\prime}}$
or
$q\leq P,\quad\frac{1}{qQ}<|\lambda|\leq\frac{1}{qQ^{\prime}}.$
In either case, there holds
$N^{\frac{1}{6}(1-\delta)}\ll\sqrt{\min\left(P,\frac{N}{Q}\right)}\ll\sqrt{q\left(1+|\lambda|N\right)}\ll\sqrt{P^{\prime}+\frac{N}{Q^{\prime}}}\ll N^{1/6}.$
Then Lemma 5.3 gives
$\displaystyle\max_{\alpha\in\mathfrak{m}\backslash\mathfrak{n}}\left|H_{2,2}(N,\alpha)\right|$$\displaystyle\ll\left\{N^{1/4}\sqrt{q\left(1+|\lambda|N\right)}+N^{2/5}+\frac{N^{1/2}}{\sqrt{q\left(1+|\lambda|N\right)}}\right\}d(q)^{2}(\log N)^{c^{\prime\prime}}$
$\displaystyle\ll N^{5/12+\varepsilon}+N^{\frac{1}{3}+\frac{\delta}{6}+\varepsilon}.$ (5.18)
Combing (5.17) and (5.3), we derive that
$\displaystyle\max_{\alpha\in\mathfrak{m}}\left|H_{2,2}(N,\alpha)\right|\ll N^{\frac{7}{16}+\varepsilon}+N^{\frac{1}{3}+\frac{\delta}{6}+\varepsilon}.$ (5.19)
It follows from (5.19), (5.2.2), (5.3) and (5.1) that
$w_{\pi}(N)\ll m^{2}(N^{\frac{23}{16}+\varepsilon}+N^{\frac{4}{3}+\frac{\delta}{6}+\varepsilon})(\log(N\mathfrak{q}_{\pi}))^{3}.$
This finishes the proof of Theorem 1.2.
Remark.
For $w_{\pi}(N)$ under Hypothesis S, if we refrain from subdividing the minor arcs further and instead individually apply Lemma 5.1 or Lemma 5.3, we obtain the following two results.
  1. (1)
    Using the second inequality of Lemma 5.1 and choosing $P=N^{\frac{2}{5}(1-\delta)}$ , we have
    $\max_{\alpha\in\mathfrak{m}}\left|H_{2,2}(N,\alpha)\right|\ll N^{1/2+\varepsilon}\left(\frac{1}{P}+\frac{1}{N^{1/4}}\right)^{1/4}\ll N^{2/5+\delta/10+\varepsilon}.$
    Inserting this into (5.2.2) and combing (5.3), we obtain
    $w_{\pi}(N)\ll m^{2}N^{\frac{7}{5}+\frac{\delta}{10}+\varepsilon}(\log(N\mathfrak{q}_{\pi}))^{3}.$
  2. (2)
    Similarly, using Lemma 5.3 and choosing $P=N^{1/2-\delta/3}$, we have
    $\max_{\alpha\in\mathfrak{m}}\left|H_{2,2}(N,\alpha)\right|\ll N^{1/2+\varepsilon}\left(N^{1/4}P^{-1/2}+N^{-1/10}\right)\ll N^{1/2+\delta/6+\varepsilon}.$
    Then
    $w_{\pi}(N)\ll m^{2}N^{3/2+\delta/6+\varepsilon}(\log(N\mathfrak{q}_{\pi}))^{3}.$
By directly comparing the aforementioned two results, it becomes evident that $N^{\frac{7}{5}+\frac{\delta}{10}+\varepsilon}$ and $N^{3/2+\delta/6+\varepsilon}$ are larger than $(N^{\frac{23}{16}+\varepsilon}+N^{\frac{4}{3}+\frac{\delta}{6}+\varepsilon})$. Hence, the rationale behind further subdividing the minor arcs is better.
5.4. Proof of Theorem 1.3
Similar to the proof of Theorem 1.1, we have
$\displaystyle\sum_{p_{1},p_{2},\dots,p_{u}\in\mathcal{J}_{k,u}(N)}(\lambda_{\pi\times\tilde{\pi}}(p_{1})-1)$ $\displaystyle=\int_{0}^{1}T^{\prime}_{\pi\times\tilde{\pi},k}(N,\alpha)\prod_{u^{\prime}=2}^{u}H_{k,u^{\prime}}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
$\displaystyle=\int_{\mathfrak{M}}T^{\prime}_{\pi\times\tilde{\pi},k}(N,\alpha)\prod_{u^{\prime}=2}^{u}H_{k,u^{\prime}}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha$
$\displaystyle\quad+\int_{\mathfrak{m}}T^{\prime}_{\pi\times\tilde{\pi},k}(N,\alpha)\prod_{u^{\prime}=2}^{u}H_{k,u^{\prime}}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha,$ (5.20)
where $H_{k,u^{\prime}}(N,\alpha)$ is defined in (5.2), $\mathfrak{M}$ and $\mathfrak{m}$ are defined in (5.1) and
$T^{\prime}_{\pi\times\tilde{\pi},k}(N,\alpha)=\sum_{p_{1}\leq N^{1/k}}(\lambda_{\pi\times\tilde{\pi}}(p_{1})-1)e(p_{1}^{k}\alpha).$
Let $k=1,u=3$ in (5.4), we use Cauchy–Schwarz inequality to get
$\displaystyle\sum_{p_{1},p_{2},p_{3}\in\mathcal{J}_{1,3}(N)}(\lambda_{\pi\times\tilde{\pi}}(p_{1})-1)$
$\displaystyle\ll\left(\max_{\alpha\in\mathfrak{M}}\left|T^{\prime}_{\pi\times\tilde{\pi},1}(N,\alpha)\right|\right)\left\{\int_{0}^{1}\left|H_{1,2}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}\left\{\int_{0}^{1}\left|H_{1,3}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}$
$\displaystyle\quad+\left(\max_{\alpha\in\mathfrak{m}}\left|H_{1,2}(N,\alpha)\right|\right)\left\{\int_{0}^{1}\left|T^{\prime}_{\pi\times\tilde{\pi},1}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}\left\{\int_{0}^{1}\left|H_{1,3}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}.$ (5.21)
Choosing $P=(\log N)^{3B+8}$, it follows from partial summation, Proposition 2.3, Hypothesis A and (5.5) that the first term above on the right side is
$\displaystyle\ll P^{1+(m^{2}+1)/2}N\exp(-c(\log N)^{1/4})N/\log N\ll N^{2}(\log N)^{-B}.$ (5.22)
By Lemma 5.1 and Hypothesis A, the second term on the right side of (5.4) is
$\displaystyle\ll NP^{-1/2}(\log N)^{3}(Nm^{4}\log^{2}(N\mathfrak{q}_{\pi\times\tilde{\pi}}))^{1/2}(N/\log N)^{1/2}\ll N^{2}(\log N)^{-B}.$ (5.23)
Inserting (5.22) and (5.23) into (5.4), we get
$r_{\pi\times\tilde{\pi}}(N)=\#\mathcal{J}_{1,3}(N)+O_{\pi,B}(N^{2}(\log N)^{-B}).$
Similarly, let $k=2,u=5$ in (5.4), then
$\displaystyle\sum_{p_{1},p_{2},p_{3},p_{4},p_{5}\in\mathcal{J}_{2,5}(N)}(\lambda_{\pi\times\tilde{\pi}}(p_{1})-1)$
$\displaystyle\ll\left(\max_{\alpha\in\mathfrak{M}}\left|T^{\prime}_{\pi\times\tilde{\pi},2}(N,\alpha)\right|\right)\left\{\int_{0}^{1}\left|H_{2,2}(N,\alpha)H_{2,3}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}\left\{\int_{0}^{1}\left|H_{2,4}(N,\alpha)H_{2,5}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}$
$\displaystyle\quad+\left(\max_{\alpha\in\mathfrak{m}}\left|H_{2,2}(N,\alpha)\right|\right)\left\{\int_{0}^{1}\left|T^{\prime}_{\pi\times\tilde{\pi},2}(N,\alpha)H_{2,3}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}\left\{\int_{0}^{1}\left|H_{2,4}(N,\alpha)H_{2,5}(N,\alpha)\right|^{2}\mathrm{d}\alpha\right\}^{1/2}.$ (5.24)
Choosing $P=\log^{9B+8c^{\prime}+16}N$, by partial summation, Proposition 2.3 and (5.7), the first term on the right side of (5.4) is
$\displaystyle\ll P^{1+(m^{2}+1)/2}N^{1/2}\exp(-c(\log N)^{1/4})N\log N\ll N^{3/2}(\log N)^{-B}.$ (5.25)
By Lemma 5.2 and Hypothesis A, the second term on the right side of (5.4) is
$\displaystyle\ll N^{1/2}(\log N)^{c^{\prime}}P^{-1/8}(Nm^{4}\log^{2}(N\mathfrak{q}_{\pi\times\tilde{\pi}})\log N)^{1/2}(N\log N)^{1/2}\ll N^{3/2}(\log N)^{-B}.$ (5.26)
Inserting (5.25) and (5.26) into (5.4), then
$w_{\pi\times\tilde{\pi}}(N)=\#\mathcal{J}_{2,5}(N)+O_{\pi,B}(N^{2}(\log N)^{-B}).$
6. Preliminaries and proof of Theorem 1.4
6.1. The modular $L$-functions
Fix a non-CM form $f\in H_{k_{1}}^{*}(N_{1})$. Let
$f(z)=\sum_{n=1}^{\infty}\lambda_{f}(n)n^{(k_{1}-1)/2}e(nz)$
be its normalized Fourier expansion at the cusp $\infty$. For each prime $p$, let $\theta_{p}\in[0,\pi]$ be the unique angle such that $\lambda_{f}(p)=2\cos\theta_{p}$. The modular $L$-function $L(s,f)$ associated with $f$ has the Euler product representation
$L(s,f)=\sum_{n=1}^{\infty}\frac{\lambda_{f}(n)}{n^{s}}=\prod_{p}\left(1-\frac{\lambda_{f}(p)}{p^{s}}+\frac{\chi_{0}(p)}{p^{2s}}\right)^{-1},\quad\mathrm{Re}(s)>1,$
where $\chi_{0}$ is the trivial Dirichlet character modulo $N_{1}$. Let’s rewrite the Euler product as
$L(s,f)=\prod_{p\mid q}\left(1-\frac{\left(-\lambda_{p}p^{-\frac{1}{2}}\right)}{p^{s}}\right)\prod_{p\nmid q}\prod_{j=0}^{1}\left(1-\frac{e^{i(2j-1)\theta_{p}}}{p^{s}}\right)^{-1},\quad\mathrm{Re}(s)>1,$
where $\lambda_{p}\in\{-1,1\}$ is the eigenvalue of the Atkin-Lehner operator $|_{k_{1}}W\left(Q_{p}\right)$. According to Deligne [7], for any prime $p$, there are two numbers $\alpha_{f}(p)$ and $\beta_{f}(p)$ such that
$\alpha_{f}(p)=e^{i\theta_{p}},\quad\beta_{f}(p)=e^{-i\theta_{p}}$ (6.1)
and
$\lambda_{f}(p)=\alpha_{f}(p)+\beta_{f}(p)$
with $\theta_{p}\in[0,\pi]$. Deligne proved the Ramanujan-Petersson conjecture
$\left|\lambda_{f}(n)\right|\leq d(n).$
For each $j\geq 1$ and $\mathrm{Re}(s)>1$, the $j$th symmetric power $L$-function attached to $f$ is defined by
$L\left(s,\mathrm{sym}^{j}f\right)=\sum_{n=1}^{\infty}\frac{\lambda_{\mathrm{sym}^{j}f}(n)}{n^{s}}=\prod_{p\mid q}\prod_{m=0}^{j}\left(1-\frac{\alpha_{\mathrm{sym}^{j}f,m}(p)}{p^{s}}\right)^{-1}\prod_{p\nmid q}\prod_{m=0}^{j}\left(1-\frac{e^{i(2m-j)\theta_{p}}}{p^{s}}\right)^{-1}.$ (6.2)
The values $\alpha_{\mathrm{sym}^{j}f,m}(p)$ can be determined by using [33, Appendix] and [6, Appendix]. Note that $L\left(s,\mathrm{sym}^{1}f\right)=L(s,f)$ and $L\left(s,\mathrm{sym}^{0}f\right)=\zeta(s)$. It is well known that $\lambda_{\mathrm{sym}^{j}f}(n)$ is a real multiplicative function. Via (6.2), one easily verify that
$\lambda_{\mathrm{sym}^{j}f}(p)=\frac{\sin(n+1)\theta_{p}}{\sin\theta_{p}}=U_{j}\left(\cos\theta_{p}\right),\quad p\nmid N_{1}.$
Let $\Gamma_{\mathbb{R}}(s):=\pi^{-s/2}\Gamma(s/2)$ and $\Gamma_{\mathbb{C}}(s)=\Gamma_{\mathbb{R}}(s)\Gamma_{\mathbb{R}}(s+1)$. With these definitions, we obtain that
$L\left(s,\left(\mathrm{sym}^{j}f\right)_{\infty}\right)=\begin{cases}q_{\mathrm{sym}^{j}f}\prod_{m=1}^{(j+1)/2}\Gamma_{\mathbb{C}}\left(s+\left(m-\frac{1}{2}\right)(k-1)\right),&\text{ if }j\text{ is odd},\\ q_{\mathrm{sym}^{j}f}\Gamma_{\mathbb{R}}(s+r)\prod_{m=1}^{j/2}\Gamma_{\mathbb{C}}(s+m(k-1)),&\text{ if }j\text{ is even},\end{cases}$ (6.3)
[…]
For all primitive Dirichlet characters $\chi$, we give a proof that the twist $L$-functions $L(s,\mathrm{sym}^{j}f\otimes\chi)$ have no Landau–Siegel zeros.
Lemma 6.1.
Let $\chi$ be a primitive Dirichlet character modulo $q$ with $q\leq Q$, $Q$ be a large positive number. There exists a constant $c_{9}>0$ such that if $n\geq 1$, then $L\left(s,\operatorname{sym}^{n}f\otimes\chi\right)\neq 0$ for
$\mathrm{Re}(s)\geq 1-\frac{c_{9}}{n^{2}\log(nk_{1}N_{1}Q(3+|\operatorname{Im}(s)|))}.$
Proof.
If $\mathrm{sym}^{n}f=\mathrm{sym}^{n}f\otimes\chi$, by [36, Corollary 4.2] and [36, Lemma 6.3], we obtain that $L(s,\mathrm{sym}^{n}f\otimes\chi)\neq 0$ for
$\mathrm{Re}(s)\geq 1-\frac{c_{9}}{n^{2}\log(nk_{1}N_{1}Q(3+|\operatorname{Im}(s)|))}.$
If $\mathrm{sym}^{n}f\neq\mathrm{sym}^{n}f\otimes\chi$, we consider the isobaric automorphic representation $\Pi=$ $\mathrm{1}\boxplus\mathrm{sym}^{2}\pi_{f}\boxplus\mathrm{sym}^{n}\pi_{f}$ Define
$(\Pi\times\widetilde{\Pi})_{\chi}:=$ $\chi\boxplus\mathrm{sym}^{2}\pi_{f}\otimes\chi\boxplus\widetilde{\mathrm{sym}^{2}\pi}_{f}\otimes\chi\boxplus\mathrm{sym}^{n}\pi_{f}\otimes\chi\boxplus\widetilde{\mathrm{sym}^{n}\pi}_{f}\otimes\chi$
$\boxplus\mathrm{sym}^{2}\pi_{f}\times\widetilde{\mathrm{sym}^{2}\pi}_{f}\otimes\chi\boxplus\widetilde{\mathrm{sym}^{2}\pi_{f}}\times\mathrm{sym}^{n}\pi_{f}\otimes\chi$
$\boxplus\mathrm{sym}^{2}\pi_{f}\times\widetilde{\mathrm{sym}^{n}\pi}_{f}\otimes\chi\boxplus\mathrm{sym}^{n}\pi_{f}\times\widetilde{\mathrm{sym}^{n}\pi}_{f}\otimes\chi,$
$\Pi^{\star}:=(\Pi\times\widetilde{\Pi})\boxplus(\Pi\times\widetilde{\Pi})_{\chi}.$
Using (1.5) and $\mathrm{sym}^{n}\pi_{f}=\widetilde{\mathrm{sym}^{n}\pi_{f}}$, $\operatorname{sym}^{n}\pi_{f}\otimes\operatorname{sym}^{2}\pi_{f}=\boxplus_{j=0}^{2}\operatorname{sym}^{n+2-2j}\pi_{f}$, $\mathrm{sym}^{n}\pi_{f}\otimes\mathrm{sym}^{n}\pi_{f}=\mathrm{1}\boxplus\left(\boxplus_{m=1}^{n}\mathrm{sym}^{2m}\pi_{f}\right)$ for $n\geq 2$, we find that for $n\geq 1$, $L\left(s,\Pi^{\star}\right)$ equals
$\zeta(s)L(s,\chi)L^{2}(s,\mathrm{sym}^{2}\pi_{f})L^{2}(s,{\mathrm{sym}^{n}\pi_{f}})L^{2}(s,\mathrm{sym}^{2}\pi_{f}\otimes\chi)L^{2}(s,{\mathrm{sym}^{n}\pi_{f}}\otimes\chi)$
$\cdot(\Pi_{m=0}^{2}L(s,\mathrm{sym}^{2m}\pi_{f}))(\Pi_{j=0}^{2}L^{2}(s,\mathrm{sym}^{n+2-2j}\pi_{f}))(\Pi_{m=0}^{n}L(s,\mathrm{sym}^{2m}\pi_{f}))$
$(\Pi_{m=0}^{2}L(s,\mathrm{sym}^{2m}\pi_{f}\otimes\chi))(\Pi_{j=0}^{2}L^{2}(s,\mathrm{sym}^{n+2-2j}\pi_{f}\otimes\chi))(\Pi_{m=0}^{n}L(s,\mathrm{sym}^{2m}\pi_{f}\otimes\chi)).$
Note that $L\left(s,\Pi^{\star}\right)$ always has a pole of order exactly 3, but a proposed zero of $L\left(s,\operatorname{sym}^{n}\pi_{f}\otimes\chi\right)$ ensures that $L\left(s,\Pi^{\star}\right)$ has a real zero of order at least 4 in the region
$\mathrm{Re}(s)\geq 1-\frac{c_{9}}{n^{2}\log(nk_{1}N_{1}Q(3+|\operatorname{Im}(s)|))}.$
This contradicts [20, Lemma 5.9]; hence, no such zero can exist (once $c_{9}$ is small enough).
6.2. The Sato–Tate conjecture
Serre’s modular version of the Sato–Tate conjecture posits that if $f$ is non-CM, then the sequence $\left\{\theta_{p}\right\}$ is uniformly distributed in the interval $[0,\pi]$ with respect to the measure $\mathrm{d}\mu_{ST}:=(2/\pi)\sin^{2}\theta\mathrm{d}\theta$. More explicitly, let $I$ be a subinterval of $[-2,2]$, and for a positive real number $x$, define the set
$N_{I}(f,x):=\#\left\{p\leq x:\mathrm{gcd}(p,N_{1})=1,\lambda_{f}(p)\in I\right\}.$
The Sato–Tate conjecture states that for a fixed non-CM form $f\in H_{k_{1}}^{*}(N_{1})$, we have
$\lim_{x\rightarrow\infty}\frac{N_{I}(f,x)}{\pi(x)}=\int_{I}\mu_{\infty}(t)\mathrm{d}t,$
where
$\mu_{\infty}(t):=\begin{cases}\frac{1}{\pi}\sqrt{1-\frac{t^{2}}{4}},&\text{ if }t\in[-2,2],\\ 0,&\text{ otherwise. }\end{cases}$
If $\chi_{I}$ is the indicator function of the interval $I=[\alpha,\beta]\subseteq[0,\pi]$, then
$\sum_{\begin{subarray}{c}p\leq x\\ \lambda_{f}(p)\in[2\cos\beta,2\cos\alpha]\end{subarray}}1=\sum_{p\leq x}\chi_{I}\left(\theta_{p}\right).$
Thorner [35] approximate $\chi_{I}$ with a differentiable function using the following construction formula [40].
Lemma 6.2.
Let $R$ be a positive integer, and let $a,b,\tau\in\mathbb{R}$ satisfy
$0<\delta^{\prime}<1/2,\quad\delta^{\prime}\leq b-a\leq 1-\delta^{\prime}.$
Then there exists an even periodic function $g(y)$ with period 1 satisfying
  1. (1)
    $g(y)=1$ when $y\in\left[a+\frac{\delta^{\prime}}{2},b-\frac{\delta^{\prime}}{2}\right]$,
  2. (2)
    $g(y)=0$ when $y\in\left[b+\frac{\delta^{\prime}}{2},1+a-\frac{\delta^{\prime}}{2}\right]$,
  3. (3)
    $0\leq g(y)\leq 1$ when $y$ is in the rest of the interval $\left[a-\frac{\delta^{\prime}}{2},1+a-\frac{\delta^{\prime}}{2}\right]$, and $g(y)$ has the Fourier expansion
    $g(y)=b-a+\sum_{n=1}^{\infty}\left(a_{n}\cos(2\pi ny)+b_{n}\sin(2\pi ny)\right),$
    where for all $n\geq 1$,
    $\left|a_{n}\right|,\left|b_{n}\right|\leq\min\left\{2(b-a),\frac{2}{n\pi},\frac{2}{n\pi}\left(\frac{R}{\pi n\delta^{\prime}}\right)^{R}\right\}.$
6.3. Proof of Theorem 1.4
If $\chi_{I}$ is the indicator function of the interval $I=[\alpha,\beta]\subseteq[0,\pi]$, then
#$\#\mathcal{J}_{1,3,f,I}(N)=\sum_{\begin{subarray}{c}(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)\\ \lambda_{f}(p_{1})\in[2\cos\beta,2\cos\alpha]\end{subarray}}1=\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\chi_{I}\left(\theta_{p_{1}}\right).$
Let $g(\theta_{p_{1}})$ be defined as in Lemma 6.2 with $a=\frac{\alpha}{2\pi}-\frac{\delta^{\prime}}{2}$ and $b=\frac{\beta}{2\pi}+\frac{\delta^{\prime}}{2}$. Thorner [35] constructed $g^{+}(\theta_{p_{1}};I,\delta^{\prime})=g\Big{(}\frac{\theta_{p_{1}}}{2\pi}\Big{)}+g\Big{(}-\frac{\theta_{p_{1}}}{2\pi}\Big{)}$ as a pointwise upper bound for $\chi_{I}(\theta_{p_{1}})$. By repeating this construction with $a=\frac{\alpha}{2\pi}+\frac{\delta^{\prime}}{2}$, and $b=\frac{\beta}{2\pi}-\frac{\delta^{\prime}}{2}$, we can get a lower bound $g^{-}(\theta_{p_{1}};I,\delta^{\prime})$ for $\chi_{I}(\theta_{p_{1}})$. Thus, according to Lemma 6.2, we can express $g^{\pm}(\theta_{p_{1}};I,\delta^{\prime})$ in terms of the basis of Chebyshev polynomials of the second kind $\{U_{n}(\cos(\theta_{p_{1}}))\}_{n=0}^{\infty}$ as follows:
$g^{\pm}(\theta_{p_{1}};I,\delta^{\prime})=a_{0}^{\pm}(I,\delta^{\prime})-a_{2}^{\pm}(I,\delta^{\prime})+\sum_{n=1}^{\infty}\left(a_{n}^{\pm}(I,\delta^{\prime})-a_{n+2}^{\pm}(I,\delta^{\prime})\right)U_{n}(\cos(\theta_{p_{1}})),$ (6.4)
where $a_{n}^{\pm}(I,\delta^{\prime})$ represents the $n$-th Fourier coefficient in the cosine expansion of $g^{\pm}(\theta_{p_{1}};I,\delta^{\prime})$. Furthermore, we derive that
$\left|a_{0}^{\pm}(I,\delta^{\prime})-a_{2}^{\pm}(I,\delta^{\prime})-\mu_{ST}(I)\right|$ $\ll\delta^{\prime},$ (6.5)
and
$\left|a_{n}^{\pm}(I,\delta^{\prime})\right|,\left|a_{n+2}^{\pm}(I,\delta^{\prime})\right|\leq\min\left\{2(b-a),\frac{2}{n\pi},\frac{2}{n\pi}\left(\frac{R}{\pi n\delta^{\prime}}\right)^{R}\right\}.$ (6.6)
Following the method in [34], when summing $g^{\pm}\left(\theta_{p_{1}};I,\delta^{\prime}\right)$ over $\mathcal{J}_{1,3}(N)$, we can interchange the order of summation by choosing $R$ to ensure absolute convergence. Thus by (6.4), we obtain the following expression:
$\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}g^{\pm}(\theta_{p_{1}};I,\delta^{\prime})$ $=\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}(a_{0}^{\pm}(I,\delta^{\prime})-a_{2}^{\pm}(I,\delta^{\prime}))$
$\quad+\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\bigg{(}\sum_{n=1}^{\infty}\left(a_{n}^{\pm}(I,\delta^{\prime})-a_{n+2}^{\pm}(I,\delta^{\prime})\right)U_{n}(\cos(\theta_{p_{1}}))\bigg{)}$
$=(a_{0}^{\pm}(I,\delta^{\prime})-a_{2}^{\pm}(I,\delta^{\prime}))\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}1$
$\quad+\sum_{n=1}^{\infty}\left(a_{n}^{\pm}(I,\delta^{\prime})-a_{n+2}^{\pm}(I,\delta^{\prime})\right)\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}U_{n}(\cos(\theta_{p_{1}})).$
It follows from (6.5), (6.6) and the above inequality that
$\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\chi_{I}\left(\theta_{p_{1}}\right)=(\mu_{\mathrm{ST}}(I)+O(\delta^{\prime}))\bigg{(}\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}1\bigg{)}+O(A),$ (6.7)
where
$\displaystyle A=\sum_{n=\delta^{\prime-1}}^{\infty}\frac{4}{n\pi}\left(\frac{R}{\pi n\delta^{\prime}}\right)^{R}\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\lambda_{\mathrm{sym}^{n}f}(p_{1})+\sum_{n=1}^{\delta^{{}^{\prime}-1}}\frac{2}{n\pi}\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\lambda_{\mathrm{sym}^{n}f}(p_{1}).$ (6.8)
Let $\pi_{f}\in\mathfrak{E}_{2}$ correspond to $f$. It is known from [28] and [29] that $\mathrm{sym}^{j}f$ corresponds to $\mathrm{sym}^{j}\pi_{f}\in\mathfrak{E}_{j+1}$. Utilizing the method outlined in the proof of Theorem 1.2, (3.7) with the zero-free region in Lemma 6.1, and [36, (6.5)], we derive the inequality:
$\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\lambda_{\mathrm{sym}^{n}f}(p_{1})\ll(n+1)N^{2}\exp\bigg{(}-\frac{c_{9}\log N}{n^{2}\log(nk_{1}N_{1}\log N)}\bigg{)}.$
Based on Lemma 6.2, we choose $\delta^{\prime-1}=(\log N)^{1/2}/\log(k_{1}N_{1}\log N)$, and $R=2$ in (6.8), then
$A\ll\frac{N^{2}(\log N)^{1/2}}{\log(k_{1}N_{1}\log N)}\exp\bigg{(}-\frac{c_{9}\log^{2}(k_{1}N_{1}\log N)}{\log(k_{1}N_{1}\log^{2}N)}\bigg{)}.$
Inserting this upper bound of $A$ into (6.7), we have
$\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\chi_{I}\left(\theta_{p_{1}}\right)=\mu_{\mathrm{ST}}(I)\left(\#\mathcal{J}_{1,3}(N)\right)+O\left(\frac{\log(k_{1}N_{1}\log N)}{(\log N)^{1/2}}\left(\#\mathcal{J}_{1,3}(N)\right)\right).$
Under Hypothesis S and utilizing the method outlined in the proof of Theorem 1.2, (3.7), and [36, (6.5)], we derive the inequality:
$\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\lambda_{\mathrm{sym}^{n}f}(p_{1})\ll(n+1)N^{5/3+\delta/3}(\log N)^{3}\log(k_{1}N_{1}n).$
In (6.8), we choose $\delta^{\prime}=RN^{\delta/6-1/6}\log(k_{1}N_{1}N)$ and $R=2$ based on Lemma 6.2, then
$A\ll N^{11/6+\delta/6}(\log N)^{3}.$
Substituting this upper bound of $A$ into (6.7), we have
$\sum_{(p_{1},p_{2},p_{3})\in\mathcal{J}_{1,3}(N)}\chi_{I}\left(\theta_{p_{1}}\right)=\mu_{\mathrm{ST}}(I)\left(\#\mathcal{J}_{1,3}(N)\right)+O(N^{11/6+\delta/6}(\log N)^{3}).$
7. Appendix
In this section, for a better understanding of the results in Theorem 1.3, we provide asymptotic formulas for the ternary Goldbach conjecture, as well as an asymptotic formula for the number of the sums of five square primes.
Lemma 7.1.
For a large integer $N$, we have
  1. (1)
    #$\#\mathcal{J}_{1,3}(N)=\mathfrak{G}_{1,3}(N)\frac{N^{2}}{2(\log N)^{3}}+O\left(\frac{N^{2}\log\log N}{(\log N)^{4}}\right),$
    where $\mathfrak{G}_{1,3}(N)$ is the singular series defined in (1.1) and satisfies $\mathfrak{G}_{1,3}(N)\geq 1/2$ for odd $N$.
  2. (2)
    #$\#\mathcal{J}_{2,5}(N)=\mathfrak{G}_{2,5}(N)\frac{\pi^{2}N^{3/2}}{24(\log N)^{5}}+O\left(\frac{N^{3/2}\log\log N}{(\log N)^{6}}\right),$
[…]
Proof.
One can see the detail proof in [17, Theorem 11]. Here we take the first equation as an example to illustrate the main ideas. By the Hardy–Littlewood circle method, we consider
#$\displaystyle\#\mathcal{J}_{1,3}(N)$ =$\displaystyle=\left\{\int_{\mathfrak{M}}+\int_{\mathfrak{m}}\right\}H_{1,1}(N,\alpha)H_{1,2}(N,\alpha)H_{1,3}(N,\alpha)e(-N\alpha)\mathrm{d}\alpha,$
where $\mathfrak{M}$ and $\mathfrak{m}$ are defined in (5.1). For the integral on the minor arcs, we follow a similar approach as in (5.2.2), but replace $T_{1,1}(N,\alpha)$ with $H_{1,1}(N,\alpha)$. Regarding the major arcs, due to the sparsity of prime numbers, we seek an equation that establishes a connection between prime numbers and integers to derive an asymptotic formula. Consider
$\displaystyle H_{1,i}(N,\alpha)=$ $\displaystyle\sum_{\begin{subarray}{c}h=1\\ (h,q)=1\end{subarray}}^{q}e\left(\frac{ah}{q}\right)\sum_{\begin{subarray}{c}\sqrt{N}<p\leq N\\ p\equiv h(\bmod q)\end{subarray}}e\left(p\lambda\right)+O(\sqrt{N})$
$\displaystyle$== $\displaystyle\mu(q)\int_{\sqrt{N}}^{N}e(u\lambda)\mathrm{d}\sum_{p\leq u,~{}p\equiv h~{}({\rm mod}~{}q)}1+O(\sqrt{N}).$
When $q\leq\log^{A}u$, by Siegel–Walfisz theorem, we have
$\sum_{p\leq u,~{}p\equiv h~{}({\rm mod}~{}q)}1=\frac{Li(u)}{\varphi(q)}+O(ue^{-c\sqrt{\log u}}).$
Moreover, under the generalized Riemann hypothesis (GRH in brief), the error term can be improved to
$O(u^{1/2}\log^{2}u).$
By substituting either of the above equations into the major arcs, and using the properties of characteristic functions with the integral in [17, Theorem 11], we get the first equation.
The proof shows the error term in the asymptotic formula cannot be improved under GRH, which is different from Theorem 1.2. One needs to explore other methods for improving the error term. ∎
Data availability statement Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Declaration
Conflict of interest On behalf of all authors, the corresponding author states that there is no conflict of interest.